### v1

# Observer-robust energy condition verification for warp drive spacetimes

Abstract
We present warpax, an open-source, GPU-accelerated Python toolkit for observer-robust energy condition analysis of warp drive spacetimes. Existing tools evaluate energy conditions for a finite sample of observer directions; warpax replaces discrete sampling with continuous, gradient-based optimization over the timelike observer manifold (rapidity and boost direction), backed by Hawking–Ellis algebraic classification. At Type I stress-energy points, which comprise > 96{>}\,96% of all grid points across the tested metrics, an algebraic eigenvalue check determines energy-condition satisfaction exactly, independent of any observer search or rapidity cap. At non-Type I points the optimizer provides rapidity-capped diagnostics. Stress-energy tensors are computed from the ADM metric via forward-mode automatic differentiation, eliminating finite-difference truncation error. Geodesic integration with tidal-force and blueshift analysis is also included.
We analyze five warp drive metrics (Alcubierre, Lentz, Van Den Broeck, Natário, Rodal) and one warp shell metric (used primarily as a numerical stress test). For the Rodal metric, the standard Eulerian-frame analysis misses violations at over $28\%$ of grid points (dominant energy condition) and over $15\%$ (weak energy condition). Even where the Eulerian frame identifies the correct violation set, observer optimization reveals that violation severity can be orders of magnitude larger (e.g. Alcubierre weak energy condition:  ${\sim}\,90{,}000\times$ at rapidity cap $\zeta_{\max}=5$, scaling as $e^{2\zeta_{\max}}$). These results demonstrate that single-frame evaluation can systematically underestimate both the spatial extent and the magnitude of energy condition violations in warp drive spacetimes. warpax is freely available at https://github.com/anindex/warpax.
keywords:
warp drive, energy conditions, Hawking–Ellis classification, general relativity, automatic differentiation, JAX
PAPER
1 Introduction
Alcubierre’s 1994 warp drive metric [1] demonstrated that general relativity admits solutions describing effective superluminal travel by deforming the spacetime geometry around a compact region, a “warp bubble”. Inside the bubble, an observer follows a timelike worldline in locally flat spacetime while the bubble itself moves at an arbitrary coordinate velocity $v_{s}$ through the external Minkowski background. The metric takes the ADM $3+1$ form
$\mathrm{d}s^{2}=-\mathrm{d}t^{2}+(\mathrm{d}x-v_{s}f(r_{s})\,\mathrm{d}t)^{2}+\mathrm{d}y^{2}+\mathrm{d}z^{2}\,$, (1)
where $f(r_{s})$ is a top-hat shaping function that smoothly transitions from unity inside the bubble to zero outside, and $r_{s}=[(x-x_{s}(t))^{2}+y^{2}+z^{2}]^{1/2}$ is the distance from the bubble center.
The fundamental obstacle to warp drive realizability is the requirement for exotic matter: matter whose stress-energy tensor violates the classical energy conditions. Throughout this paper, the spacetime signature is (+++)(-+++) and we work in geometric units ($G=c=1$). The four energy conditions, namely the null (NEC), weak (WEC), strong (SEC), and dominant (DEC), are pointwise inequalities on the stress-energy tensor contracted with timelike or null vectors:
NEC: $\quad T_{ab}\,k^{a}k^{b}\geq 0$ $\forall\;\text{null }k^{a}\,$, (2)
WEC: $\quad T_{ab}\,u^{a}u^{b}\geq 0$ $\forall\;\text{timelike }u^{a}\,$, (3)
SEC: $\quad\left(T_{ab}-\tfrac{1}{2}\,T\,g_{ab}\right)u^{a}u^{b}\geq 0$ $\forall\;\text{timelike }u^{a}\,$, (4)
DEC: $\quad T_{ab}\,u^{a}u^{b}\geq 0\;\text{and}\;{-T^{a}{}_{b}\,u^{b}}\;\text{future causal}$ $\forall\;\text{future timelike }u^{a}\,.$ (5)
These conditions must hold for all admissible observers at each spacetime point, not merely for a single preferred frame [2, 3]. The truth value of an energy condition at a point is observer-independent (it is defined by a universal quantifier); what is observer-dependent is any single-frame diagnostic that evaluates $T_{ab}\,u^{a}u^{b}$ for a particular $u^{a}$.
The WarpFactory toolkit [4] addresses this by evaluating energy conditions over a discrete set of sampled observer velocities and directions (typically $\sim$1000 null and timelike vectors), taking the minimum across all samples at each grid point. This sampling-based approach is a significant advance over single-frame evaluation, but has inherent limitations: the accuracy depends on the sampling density, and the discrete set may miss narrow violation cones in the observer manifold. WarpFactory also relies on fourth-order central finite differences for the curvature computation [5], which introduces truncation error and requires a step-size parameter.
The observer-dependence of energy condition evaluation is well-established in the mathematical relativity literature. Hawking and Ellis [3] classified the algebraic types of the stress-energy tensor (Types I–IV), and the conditions under which each energy condition can be violated depend on this algebraic structure. For Type I (the generic case), the NEC is satisfied for all observers if and only if $\rho+p_{i}\geq 0$ for each principal pressure $p_{i}$, but evaluating this requires computing the eigenvalues of the mixed tensor $T^{a}{}_{b}$, rather than simply reading off components in a coordinate basis.
In this paper, we present warpax, a JAX-based toolkit that replaces discrete observer sampling with continuous, gradient-based observer optimization over the observer manifold. warpax is built on JAX [6] for GPU-accelerated automatic differentiation, uses Equinox [7] for module structure, and provides:
  • exact derivatives of the implemented (regularized) metric functions (up to floating-point roundoff) via forward-mode automatic differentiation, eliminating finite-difference truncation error (the computed curvature is exact for the numerical profile at each point, which may differ from the target analytic metric where smoothing or clamping is applied (see Section 3),
  • Hawking–Ellis algebraic classification of the stress-energy type,
  • observer optimization over the observer manifold parameterized by rapidity to find worst-case observers at each point,
  • geodesic integration with tidal-force computation via the geodesic deviation equation, and
  • kinematic scalar analysis (expansion, shear, vorticity) for arbitrary congruences.
We analyze five warp drive metrics and one warp shell geometry. For the Rodal metric, the Eulerian single-frame DEC analysis fails to detect violations at over 28% of all grid points; for WarpShell (a regularized implementation used primarily as a numerical stress test; see Section 7.3), missed WEC violations are below $0.1\%$ of the grid. For Alcubierre and Natário, we find no additional NEC or WEC violating grid points beyond the baseline Eulerian analysis (missed =0=0 at the sampled resolution); Lentz shows zero missed NEC/WEC points but nonzero SEC misses, and is severely under-resolved ($\ll\!1$ cell across the wall at $50^{3}$); this result should be treated as a lower bound. Van Den Broeck exhibits small but nonzero missed fractions (NEC $0.1\%$, WEC $0.4\%$ at $v_{s}=0.5$). The observer-optimized violation severity (at $\zeta_{\max}=5$) can be substantially larger in regions already identified as violating. We validate against WarpFactory’s published Eulerian results and quantify the additional violations discovered by observer optimization where they exist.
Related work.
Computational tools for general relativity span a broad spectrum from symbolic computer algebra to large-scale numerical evolution (see Ref. [8] for a comprehensive survey). Symbolic packages such as xAct [9] for Mathematica and SageManifolds [10] for SageMath provide exact tensor calculus on differentiable manifolds, enabling analytic derivation of curvature tensors and energy condition expressions in closed form. These tools are invaluable for deriving exact results but do not perform numerical verification of energy conditions on discretized spacetime grids or systematic optimization over observer families. At the numerical end, community frameworks for numerical relativity (e.g. the Einstein Toolkit, NRPy+) focus on dynamical evolution of Einstein’s equations rather than pointwise energy condition assessment of static metrics. The use of automatic differentiation for computing curvature tensors in general relativity is an emerging paradigm: libraries such as diffjeom [11] demonstrate JAX-based computation of the Christoffel–Riemann–Ricci hierarchy, and Bara [12] identifies solving field equations in exotic spacetimes as an underexplored domain for gradient-based methods. To the best of our knowledge, warpax is the first tool to combine exact autodiff-derived curvature with systematic, gradient-based optimization over the timelike observer manifold for energy condition verification.
The paper is organized as follows. Section 2 presents the mathematical methods. Section 3 describes the warpax toolkit. Section 4 presents energy condition results. Section 5 presents geodesic results. Section 6 analyses kinematic scalars. Section 7 validates numerical convergence. Section 8 discusses implications, and Section 9 concludes. Appendices provide supplementary material: transition-function ablation (Appendix A), Rodal DEC ablation (Appendix B), wall-resolution evidence (Appendix C), and full metric definitions (Appendix D).
2 Methods
2.1 Stress-energy via automatic differentiation
The Einstein field equations in geometric units ($G=c=1$) are
$G_{ab}=8\pi\,T_{ab}\,$, (6)
where the Einstein tensor $G_{ab}=R_{ab}-\frac{1}{2}R\,g_{ab}$ is constructed from the Ricci tensor $R_{ab}$ and scalar curvature $R=g^{ab}R_{ab}$. Computing $T_{ab}$ from a given metric requires the full curvature chain:
$g_{ab}\;\xrightarrow{\partial}\;\Gamma^{a}{}_{bc}\;\xrightarrow{\partial}\;R^{a}{}_{bcd}\;\xrightarrow{\text{contract}}\;R_{ab},\,R\;\xrightarrow{\text{EFE}}\;T_{ab}\,.$ (7)
The Christoffel symbols of the second kind are
$\Gamma^{a}{}_{bc}=\frac{1}{2}\,g^{ad}\left(\partial_{c}\,g_{bd}+\partial_{b}\,g_{cd}-\partial_{d}\,g_{bc}\right)\,$, (8)
and the Riemann curvature tensor is
$R^{a}{}_{bcd}=\partial_{c}\,\Gamma^{a}{}_{bd}-\partial_{d}\,\Gamma^{a}{}_{bc}+\Gamma^{a}{}_{ce}\,\Gamma^{e}{}_{bd}-\Gamma^{a}{}_{de}\,\Gamma^{e}{}_{bc}\,$. (9)
warpax implements this chain using JAX’s forward-mode automatic differentiation (jax.jacfwd). The metric function $g_{ab}(x^{\mu})$ maps a coordinate 4-vector to a $4\times 4$ symmetric matrix. A single call to jacfwd yields the metric derivatives $\partial_{c}\,g_{ab}$ as exact derivatives of the implemented program, up to floating-point roundoff, with no step-size parameter and no finite-difference truncation error. All computations use 64-bit floating point (jax_enable_x64 = True). A nested application of jacfwd (forward-over-forward) computes the second derivatives needed for the Riemann tensor. For dimension $d=4$, this forward-mode approach is efficient: the Jacobian has only 4 directional derivatives, compared to the $16$ reverse-mode passes that would be needed for the $4\times 4\to 4\times 4$ Christoffel computation.
The derivative convention throughout warpax places the differentiation index last in the output array: $(\partial_{c}\,g)_{abc}$ has $\partial g_{ab}/\partial x^{c}$ at index $[a,b,c]$. This choice simplifies the einsum contractions in subsequent steps.
2.2 Hawking–Ellis classification
The algebraic classification of $T^{a}{}_{b}$ [3] determines the structure of the energy condition constraints. At each spacetime point, we compute the eigenvalues of the mixed stress-energy tensor $T^{a}{}_{b}=g^{ac}T_{cb}$ and classify the result into one of four types:
  • Type I: Four real eigenvalues with a timelike eigenvector. The generic case for physically reasonable matter. For signature (+++)(-+++), the eigenvalues of $T^{a}{}_{b}$ are $(-\rho,p_{1},p_{2},p_{3})$; we identify the timelike eigenvector via $g_{ab}\,v^{a}v^{b}<0$ and set $\rho=-\lambda_{\text{timelike}}$, $p_{i}=\lambda_{\text{spacelike},i}$.
  • Type II: A double null eigenvector (radiation-like). Stress-energy has the form of a null fluid plus a Type I part.
  • Type III: A triple null eigenvector (pure radiation with no rest frame). Rare in practice.
  • Type IV: Complex eigenvalues. Absent for most classical matter models, but possible for exotic stress–energy and also susceptible to spurious classification from numerical noise near eigenvalue degeneracies.
warpax implements a branchless classification algorithm suitable for GPU execution via jax.vmap. The classifier avoids Python-level conditionals by computing a continuous score for each type and selecting via branchless masking. The mixed tensor $T^{a}{}_{b}=g^{ac}T_{cb}$ is generally not symmetric in the Euclidean sense, so the eigendecomposition uses a general (non-symmetric) solver; eigenvalues are classified as real when $|\text{Im}\,\lambda_{i}|<\epsilon\,s$ (default $\epsilon=10^{-10}$, scale $s=\max(|\mathrm{Re}\,\lambda_{j}|,1)$). Eigenvector causal character and eigenvalue degeneracy use the same tolerance with the same relative scale factor. A second, relative tolerance tier catches split degenerate eigenvalue pairs at large $\|T^{a}{}_{b}\|$ (Section 8). Across all grids in this study, $\max|\text{Im}\,\lambda_{i}|/s<10^{-12}$ typically at Type I-classified points, confirming well-behaved eigenstructure (non-Type I points, including transition-zone Type IV, may have larger imaginary parts; see Table 8). Near-vacuum points ($\max|\text{Re}\,\lambda_{i}|<\epsilon$) are detected and assigned Type I with $\rho=p_{i}=0$ before the eigendecomposition, avoiding numerically meaningless eigenvectors. At exactly vacuum points ($T_{ab}=0$, e.g. Minkowski or Schwarzschild exterior), all eigenvalues are degenerate at zero and the timelike eigenvector is not unique; the classifier correctly labels these as Type I with $\rho=p_{i}=0$. For Type I, all energy conditions reduce to algebraic inequalities in the eigenvalues $\rho,p_{1},p_{2},p_{3}$:
NEC: $\quad\rho+p_{i}\geq 0$ $\forall\,i\,$, (10)
WEC: $\quad\rho\geq 0\;\text{and}\;\rho+p_{i}\geq 0$ $\forall\,i\,$, (11)
SEC: $\quad\rho+p_{i}\geq 0\;\;\forall\,i\;\;\text{and}\;\;\rho+{\textstyle\sum_{i}}p_{i}\geq 0\,$, (12)
DEC: $\quad\rho\geq|p_{i}|$ $\forall\,i\,.$ (13)
These inequalities are exact: for Type I stress-energy, a point satisfies (or violates) an energy condition if and only if the corresponding eigenvalue inequality holds, regardless of which observer is chosen. We therefore distinguish three diagnostic quantities:
  1. 1.
    Boolean truth (cap-independent): determined by the sign of the eigenvalue inequalities for Type I.
  2. 2.
    Algebraic slack: finite, cap-independent scalars such as $\min_{i}(\rho+p_{i})$ (NEC), $\min(\rho,\,\min_{i}(\rho+p_{i}))$ (WEC), and $\min_{i}(\rho-|p_{i}|)$ (DEC), quantifying proximity to the violation boundary.
  3. 3.
    Capped observer extremum: the $\zeta_{\max}$-dependent contraction minimum $\min_{\zeta\leq\zeta_{\max}}T_{ab}\,u^{a}u^{b}$, which for NEC-violating points diverges as $\gamma^{2}$ and whose magnitude is a diagnostic of violation severity at a chosen rapidity scale.
Throughout this paper, “robust margin” refers to the merged quantity (hereafter hybrid margin): algebraic slack at Type I points and $\zeta_{\max}$-capped extremum at non-Type I points. For Type I points, violation detection and the reported margins use the algebraic slack (item 2), which is cap-independent and authoritative. The observer optimizer is run additionally to provide the explicit worst-case observer parameters $(\zeta^{*},\theta^{*},\phi^{*})$; its $\zeta_{\max}$-capped contraction minimum (item 3) is reported separately in sensitivity analyses (Table 9) but is not used for violation detection. The algebraic slack and capped extremum are in general different numerical values: the former represents the global eigenvalue-based quantity, while the latter depends on $\zeta_{\max}$. These two quantities are conceptually distinct and should not be compared directly; the algebraic slack measures distance to the violation boundary, while the capped extremum quantifies the worst-case energy density within the tested observer range. The algebraic slack (e.g. $\min_{i}(\rho+p_{i})$ for NEC) certifies satisfaction or violation but is not, in general, equal to the minimum observer-contracted energy density $\min_{u}T_{ab}\,u^{a}u^{b}$; it measures proximity to the eigenvalue inequality boundary rather than a physical extremum over the observer manifold. For the metrics and rapidity caps used here ($\zeta_{\max}=5$) we empirically observe sign agreement between the algebraic Type I truth and the $\zeta_{\max}$-capped optimizer at all grid points (i.e. no violation is “hidden” above the cap). This is not guaranteed in general: a Type I tensor with $\rho+p_{i}<0$ but $|\rho+p_{i}|\ll|\rho|$ could require an arbitrarily large boost to flip the sign of the measured energy density, so a finite $\zeta_{\max}$ might in principle miss it. At non-Type-I grid points (up to $\sim\!3$% for the warp metrics and <1<\!1% for WarpShell; see Table 8), the algebraic inequalities do not apply and the optimizer is the sole method; this pathway is validated via a null-dust benchmark (Section 8).
2.3 Observer parameterization
To verify energy conditions over all admissible observers, we parameterize the space of timelike and null vectors at each point.
Timelike observers.
Any unit timelike vector $u^{a}$ at a point can be written as a Lorentz boost of the Eulerian normal $n^{a}$:
$u^{a}=\cosh\zeta\;n^{a}+\sinh\zeta\;\hat{s}^{a}(\theta,\phi)\,$, (14)
where $\zeta\geq 0$ is the rapidity (boost magnitude), $(\theta,\phi)$ specify the spatial boost direction, and $\hat{s}^{a}(\theta,\phi)$ is a unit spatial vector constructed from an orthonormal tetrad adapted to the ADM decomposition. The tetrad is built by applying Gram–Schmidt orthogonalization to the coordinate basis vectors projected into the spatial hypersurface orthogonal to $n^{a}$.
Null observers.
Any future-directed null vector can be written as
$k^{a}=n^{a}+\hat{s}^{a}(\theta,\phi)\,$, (15)
where $(\theta,\phi)$ parameterize the null direction on the celestial sphere. This fixes $k_{a}n^{a}=-1$, removing the null-vector rescaling ambiguity and making the NEC margin $T_{ab}\,k^{a}k^{b}$ scale-fixed. The NEC is thus a two-dimensional optimization problem over $(\theta,\phi)$.
The rapidity parameterization has the advantage that $\zeta=0$ recovers the Eulerian observer, so the optimization landscape always includes the Eulerian result as a baseline. The WEC and DEC involve three-dimensional optimization over $(\zeta,\theta,\phi)$, while the NEC requires only $(\theta,\phi)$. The true timelike observer manifold has unbounded rapidity ($\zeta\in[0,\infty)$); in practice, we optimize over $\zeta\in[0,\zeta_{\max}]$ with a fixed cap $\zeta_{\max}=5$ (corresponding to $\gamma\approx 74$; see Section 2.4).
2.4 Observer optimization
For each spacetime point and energy condition, warpax solves the optimization problem
$\text{margin}^{*}=\min_{(\zeta,\theta,\phi)\in\mathcal{O}}\;m(T_{ab},u^{a}(\zeta,\theta,\phi))\,$, (16)
where $m$ is the signed margin function: $m\geq 0$ means the condition is satisfied, and $m<0$ indicates a violation. For the NEC, $m=T_{ab}\,k^{a}k^{b}$ with optimization over $(\theta,\phi)$ only. For the DEC, the diagnostic at each spacetime point returns
$m_{\text{DEC}}=\min\bigl(m_{\text{flux}},\;m_{\text{future}},\;m_{\text{WEC}}\bigr)\,$, (17)
where $m_{\text{flux}}=-g_{ab}\,j^{a}j^{b}$ (positive when the energy flux $j^{a}=-T^{a}{}_{b}\,u^{b}$ is causal), $m_{\text{future}}=-j_{a}n^{a}$ (positive when $j^{a}$ is future-directed, with $n^{a}$ the future-pointing unit normal), and $m_{\text{WEC}}=T_{ab}\,u^{a}u^{b}$ (positive when the energy density is non-negative). A negative value of any component indicates a DEC violation for that observer. The three-term minimum in equation (17) is piecewise smooth but not $C^{1}$ at branch-switching boundaries; gradients exist almost everywhere, and the multi-start BFGS optimizer handles this reliably in practice. A log-sum-exp soft-minimum approximation could be employed if smoother gradients are needed. Because $m_{\text{WEC}}$ has dimensions of energy density while $m_{\text{flux}}$ has dimensions of (energy density)2, we do not combine them into a single margin for severity purposes. For violation detection, the DEC is violated whenever $m_{\text{DEC}}^{*}<0$ (equation 17); the sign logic is correct regardless of units. For Type I stress-energy with $\rho\geq 0$ and $\rho\geq|p_{i}|$, the algebraic condition (equation 13) guarantees that $j^{a}$ is causal and future-directed for every future-directed timelike $u^{a}$, making the three-term check redundant; the algebraic slack $\min_{i}(\rho-|p_{i}|)$ then serves as the primary DEC result with homogeneous dimensions (energy density). At non-Type I points, the optimizer evaluates all three terms of equation (17) simultaneously (Section 8). For the majority of grid points, which classify as Type I, the algebraic slack is the primary DEC result; at non-Type I points the capped extremum is used instead.
The optimization uses BFGS [13] implemented in Optimistix [14], a pure-JAX optimization library. For timelike observers (WEC/SEC/DEC), the optimizer works over an unconstrained 3-vector $\mathbf{w}\in\mathbb{R}^{3}$, with the rapidity $\zeta=|\mathbf{w}|$ and boost direction $\hat{s}=\mathbf{w}/|\mathbf{w}|$. This parameterization naturally includes the Eulerian observer at $\mathbf{w}=0$ (the removable singularity $\mathbf{w}/|\mathbf{w}|$ at the origin is handled via the smooth approximation $|\mathbf{w}|_{\varepsilon}=\sqrt{\mathbf{w}\cdot\mathbf{w}+\varepsilon^{2}}$ with $\varepsilon=10^{-12}$, which is $C^{\infty}$ everywhere including $\mathbf{w}=0$ and avoids discontinuous gradients that could confuse quasi-Newton Hessian updates near the Eulerian point; see Section 8). This parameterization avoids angular coordinate singularities and ensures a smooth optimization landscape. A soft cap $|\mathbf{w}|\mapsto\zeta_{\max}\tanh(|\mathbf{w}|/\zeta_{\max})$ bounds the rapidity at $\zeta_{\max}=5$ (corresponding to $\gamma=\cosh 5\approx 74$); the $\tanh$ map is chosen because it smoothly saturates with unit gradient near the origin, preserving optimizer sensitivity for moderate boosts while enforcing a strict upper bound. For the NEC, the null direction is parameterized via stereographic projection: a 2-vector $\mathbf{w}\in\mathbb{R}^{2}$ maps to a direction on $S^{2}$ via $\hat{s}=(2w_{1},\,2w_{2},\,1-|\mathbf{w}|^{2})/(1+|\mathbf{w}|^{2})$, then $k^{a}=e_{0}^{a}+\hat{s}^{i}\,e_{i}^{a}$. This covers all of $S^{2}$ smoothly without polar coordinate singularities ($\mathbf{w}=0\to$ north pole; $|\mathbf{w}|\to\infty\to$ south pole). Because the south pole requires $|\mathbf{w}|\to\infty$, optimizer convergence near it can be slow; the multi-start pool includes near-pole initializations (axis-aligned directions with large $|\mathbf{w}|$) to mitigate this.
Multi-start optimization is implemented via jax.vmap over $N_{\text{starts}}$ initial conditions. The first start is always the Eulerian observer ($\zeta=0$), followed by six axis-aligned spatial directions; the remaining $N_{\text{starts}}-7$ starts are drawn from a standard normal distribution. This deterministic seeding ensures the physically motivated baseline is always included. The entire multi-start pool is evaluated in parallel, which parallelises trivially on GPU hardware with no additional implementation effort.
Interpretation caveat.
Because the rapidity is capped at $\zeta_{\max}$, the reported WEC/SEC/DEC minima are $\zeta_{\max}$-capped extrema rather than true infima over all timelike observers. At spacetime points where the NEC is violated ($T_{ab}\,k^{a}k^{b}<0$ for some null $k^{a}$), the WEC energy density $T_{ab}\,u^{a}u^{b}$ diverges as $\sim\!\gamma^{2}\,T_{ab}\,k^{a}k^{b}$ for boosted observers approaching that null direction, so the true WEC infimum is -$\infty$ and only the capped value is physically meaningful. All reported WEC/SEC/DEC minima should therefore be understood as $\text{WEC}(\zeta_{\max})$ etc., that is, diagnostics conditioned on the rapidity cap rather than observer-independent scalars. For Type I points, the algebraic slacks (equations 1013) provide cap-independent Boolean violation detection. The NEC margin, which involves only the two-dimensional null direction and no rapidity, is unaffected by this caveat.
2.5 Geodesic integration
warpax integrates the geodesic equation
$\frac{\mathrm{d}^{2}x^{\mu}}{\mathrm{d}\lambda^{2}}+\Gamma^{\mu}{}_{\alpha\beta}\,\frac{\mathrm{d}x^{\alpha}}{\mathrm{d}\lambda}\,\frac{\mathrm{d}x^{\beta}}{\mathrm{d}\lambda}=0$ (18)
using the Tsitouras 5(4) Runge–Kutta method (Tsit5) [15] via Diffrax [16], with adaptive step-size control (PID controller, $\mathrm{rtol}=\mathrm{atol}=10^{-10}$).
Tidal forces are computed via the geodesic deviation equation
$\frac{D^{2}\xi^{\mu}}{\mathrm{d}\tau^{2}}=-R^{\mu}{}_{\alpha\beta\gamma}\,\frac{\mathrm{d}x^{\alpha}}{\mathrm{d}\tau}\,\xi^{\beta}\,\frac{\mathrm{d}x^{\gamma}}{\mathrm{d}\tau}\,$, (19)
which is integrated simultaneously with the geodesic equation as a coupled 16-component ODE system $y=(x^{\mu},v^{\mu},\xi^{\mu},w^{\mu})$, where $v^{\mu}=\mathrm{d}x^{\mu}/\mathrm{d}\lambda$ and $w^{\mu}=D\xi^{\mu}/\mathrm{d}\lambda$. The explicit first-order form is
$\dot{x}^{\mu}$ $=v^{\mu}\,$, $\dot{v}^{\mu}$ $=-\Gamma^{\mu}{}_{\alpha\beta}\,v^{\alpha}v^{\beta}\,$,
$\dot{\xi}^{\mu}$ $=w^{\mu}-\Gamma^{\mu}{}_{\alpha\beta}\,v^{\alpha}\xi^{\beta}\,$, $\dot{w}^{\mu}$ $=-R^{\mu}{}_{\nu\rho\sigma}\,v^{\nu}\xi^{\rho}v^{\sigma}-\Gamma^{\mu}{}_{\alpha\beta}\,w^{\alpha}v^{\beta}\,$, (20)
where dots denote $\mathrm{d}/\mathrm{d}\lambda$. The tidal tensor $K^{\mu}{}_{\rho}=R^{\mu}{}_{\nu\rho\sigma}\,v^{\nu}\,v^{\sigma}$ is a $4\times 4$ matrix with one zero eigenvalue (corresponding to the velocity direction $v^{\mu}$). The three non-zero eigenvalues coincide with those of the spatial tidal matrix $\mathcal{E}_{ij}=R_{\hat{0}i\hat{0}j}$ in an orthonormal frame comoving with the observer, and directly quantify tidal stretching and compression experienced by a freely falling observer.
Photon blueshift is computed as the ratio of the observed frequency at two points along a null geodesic:
$\frac{\nu_{\text{obs}}}{\nu_{\text{emit}}}=\frac{(g_{\mu\nu}\,u^{\mu}_{\text{obs}}\,k^{\nu})_{\text{obs}}}{(g_{\mu\nu}\,u^{\mu}_{\text{emit}}\,k^{\nu})_{\text{emit}}}\,$, (21)
where $k^{\mu}$ is the photon 4-momentum and $u^{\mu}$ is the observer 4-velocity at each endpoint.
3 The warpax toolkit
warpax is a Python package built on JAX [6] for hardware-accelerated numerical computation with automatic differentiation. The toolkit is organized into six computational modules:
  • warpax.geometry – Metric specification, Christoffel symbols, curvature tensors, stress-energy, curvature invariants, and grid evaluation with vmap-based batching.
  • warpax.energy_conditions – Hawking–Ellis classification, observer construction, observer optimization, and the unified grid verifier.
  • warpax.metrics: Six warp/shell metrics (Alcubierre [1], Lentz [17], Natário [18], Van Den Broeck [19], Rodal [20], WarpShell) and two benchmark metrics (Minkowski, Schwarzschild in isotropic Cartesian coordinates), each as an Equinox [7] module with dynamic physical parameters (Appendix D gives the full ADM decompositions).
  • warpax.geodesics: Geodesic integrator, initial condition constructors, deviation equation solver, and observable computations (proper time, blueshift, tidal eigenvalues).
  • warpax.analysis: Multi-metric comparison, Richardson extrapolation for convergence studies, and kinematic scalar computation.
  • warpax.visualization: Publication-quality figure generation: comparison panels, direction fields, convergence plots, and kinematic scalar plots.
Computational approach.
Each metric is an Equinox module implementing a pointwise signature $x^{\mu}\mapsto g_{ab}(x)$, where $x^{\mu}$ is a 4-vector of coordinates and $g_{ab}$ is a $4\times 4$ symmetric matrix. Physical parameters (bubble velocity $v_{s}$, size $R$, thickness $\sigma$) are dynamic fields that can be modified without recompilation. Grid evaluation lifts the pointwise function to arbitrary grids via jax.vmap, with optional chunked batching via jax.lax.map for memory-limited hardware.
All derivatives are computed via jax.jacfwd (forward-mode autodiff), yielding exact derivatives of the implemented metric functions up to floating-point roundoff. The full curvature chain from metric to stress-energy is JIT-compiled as a single fused computation, eliminating intermediate memory allocation. On an NVIDIA A100 GPU, evaluating the complete curvature chain on a $50^{3}$ grid takes on the order of seconds; the $100^{3}$ evaluation needed for convergence studies takes on the order of tens of seconds.
Verification procedure.
The grid verifier (warpax.energy_conditions.verifier) orchestrates the full analysis for a given metric and grid:
  1. 1.
    Evaluate the curvature chain on the grid (metric $\to$ $T_{ab}$).
  2. 2.
    Classify each point via Hawking–Ellis (Type I/II/III/IV).
  3. 3.
    For Type I points: compute algebraic slacks.
  4. 4.
    For all points: run observer optimization with multi-start (Section 2.4).
  5. 5.
    For Type I: use algebraic slacks for violation detection; report capped extrema as $\zeta_{\max}$-dependent severity diagnostics. For non-Type-I: use capped extrema. Compute per-condition summaries.
The Eulerian-only baseline evaluates the NEC for six null vectors $k^{a}=n^{a}\pm e^{a}_{(i)}$, $i=1,2,3$, built from the Eulerian orthonormal triad $\{e^{a}_{(i)}\}$ (a restricted-direction baseline; the full NEC requires optimization over all null directions), and evaluates WEC, SEC, and DEC using the single Eulerian (ADM normal) observer. Since $m^{*}(x)=\min_{\text{observers}}\,m(x,u)\leq m(x,u_{0})$ for any fixed observer $u_{0}$, the baseline margin is an upper bound on the worst-case margin at each point, and the set of baseline-detected violations is a lower bound on the true violation extent. This inequality is enforced numerically as a consistency check in the test suite.
Metric regularity.
The Alcubierre, Lentz, Natário, Rodal, and Van Den Broeck metric implementations use smooth ($C^{\infty}$) $\tanh$-based shaping functions, ensuring that all derivatives needed for the Riemann tensor are well-defined. The WarpShell metric uses $C^{2}$-smooth quintic Hermite-smoothstep transitions between shell regions. The lapse blends smoothly to unity beyond $R_{2}+\text{smooth\_width}$, the shift blends smoothly to zero beyond $R_{2}$, and the spatial metric uses a $\tanh$-based shell indicator for the conformal factor; all three components employ smooth ($C^{2}$) transition functions with no discontinuous clamping. For our parameter choice $R_{1}=0.5$, the interior is only approximately flat (the shell indicator evaluates to $\sim\!0.02$ at $r=0$). The spatial metric is only asymptotically flat, with residual curvature decaying exponentially beyond $R_{2}$. Since the curvature is computed via autodiff of the implemented functions, the computed stress-energy is the exact derivative of the numerical profile at each point; for WarpShell points near the shell boundaries, this may include artifacts from the finite-width transition rather than distributional contributions. For metrics defined via a radial distance $r_{s}=\sqrt{\cdots}$, the implementations use smooth $C^{\infty}$ regularization ($r_{s}\to\sqrt{r_{s}^{2}+\varepsilon^{2}}$ with $\varepsilon=10^{-12}$) in expressions that divide by $r_{s}$, preventing NaN propagation through the autodiff chain at the bubble center while preserving differentiability for curvature computation.
Reproducibility.
All results in this paper are generated by scripts included in the warpax repository: run_analysis.py (energy condition grid analysis at multiple velocities), run_convergence.py (Richardson extrapolation), run_geodesics.py (geodesic integration and tidal forces), run_kinematic_scalars.py (kinematic scalar evaluation), and reproduce_figures.py (figure generation from cached results). The full analysis can be reproduced with three commands.
4 Results: Energy conditions
We evaluate energy conditions on $50^{3}$ spatial grids for each metric, with the time coordinate fixed at $t=0$ (for constant-velocity bubbles whose metric depends on $x-v_{s}t$, the $t=0$ slice is representative up to spatial translation). The grid extends $\pm 5R$ for metrics with $R=1$ (Alcubierre, Van Den Broeck, Natário) and $\pm 3R$ for large-radius metrics (Rodal, Lentz with $R=100$), centered on the bubble; the WarpShell grid extends $\pm 5R_{2}$. For the Alcubierre metric with $\sigma=8$, the shape function $f(r_{s})=\{\tanh[\sigma(r_{s}+R)]-\tanh[\sigma(r_{s}-R)]\}/[2\tanh(\sigma R)]$ has a 10–90% transition width of $0.27$ coordinate units (Table 2), so the $50^{3}$ grid ($h=10/49\approx 0.20$) places $\sim\!1.3$ cells across the wall. For a monotone $\tanh$ wall $g(r)=[1-\tanh(\sigma(r-R))]/2$, the 10–90% levels correspond to $\tanh=\mp 0.8$, giving $\Delta r=2\,\mathrm{atanh}(0.8)/\sigma\approx 2.197/\sigma$. The convergence study (Section 7) confirms that the minimum NEC margin is resolution-stable (within $\sim\!1$% between $50^{3}$ and $100^{3}$) despite this marginal resolution. Table 1 lists the parameters used for each metric.
Table 1: Per-metric parameters and grid configuration. All grids are $50^{3}$ at $t=0$, centered on the bubble.
Metric $R$ $\sigma$ Domain $\zeta_{\max}$
Alcubierre 1 8.0 $(\pm 5)^{3}$ 5
Van Den Broeck 1 8.0 $(\pm 5)^{3}$ 5
Natário 1 8.0 $(\pm 5)^{3}$ 5
Rodal 100 0.03 $(\pm 300)^{3}$ 5
Lentz 100 8.0 $(\pm 300)^{3}$ 5
WarpShell $R_{1}\!=\!0.5$, $R_{2}\!=\!1$ $(\pm 5)^{3}$ 5
Schwarzschild M=1M\!=\!1 $(\pm 20)^{3}$ 5
 Additional VdB parameters: $\tilde{R}=1$, $\alpha_{\text{vdb}}=0.5$, $\sigma_{B}=8$.
 WarpShell parameters: $r_{\mathrm{S}}=5$, smooth_width =0.06=0.06, $C^{2}$ quintic Hermite transitions.
Table 2: Wall resolution analysis for all metrics ($50^{3}$ grid). Transition width computed as the 10–90% threshold distance of each metric’s shape function ($\Delta r=2\,\mathrm{atanh}(0.8)/\sigma$ for $\tanh$-based walls). Autodiff computes exact curvature at each grid point; wall resolution affects spatial sampling density, not curvature accuracy.
Metric Wall width $\Delta x$ Cells Resolved
Alcubierre 0.27 0.20 1.35 No
Van Den Broeck 0.27 0.20 1.35 No
Natário 0.27 0.20 1.35 No
Lentz 0.27 12.24 0.02 No
Rodal 73.24 12.24 5.98 Yes
WarpShell 0.06 0.20 0.29 No
We emphasize that autodiff computes exact curvature at each sampled point; under-resolution affects only the spatial sampling of the violation region, not the curvature accuracy at any given point. Nevertheless, for the most severely under-resolved case (Lentz, ${\sim}0.02$ cells across the wall), the uniform grid may fail to place any points inside the narrow transition shell, so the reported violation fractions should be interpreted as lower bounds on the true spatial extent. An adaptive or wall-clustered grid would provide tighter volume estimates for thin-wall metrics; this is deferred to future work.
For each grid point, we compute: (i) the Eulerian baseline (ADM normal observer for WEC/SEC/DEC; six axis-aligned null directions for NEC), and (ii) the observer-optimized margins via multi-start optimization (Section 2.4) with $N_{\text{starts}}=8$. The Hawking–Ellis classifier identifies the vast majority of grid points as Type I (four real eigenvalues with a timelike eigenvector): Rodal achieves 100%, Lentz 99.9%, WarpShell 99.6%, VdB 98.4%, Alcubierre 98%, and Natário $\sim\!97.5$%. The minimum across all tested metrics is Natário at 97.5% (Table 8). At Type I points the algebraic eigenvalue inequalities (equations 1013) apply; at non-Type I points the optimizer provides capped extrema as the sole diagnostic.
4.1 Individual metric analysis
Alcubierre metric.
The original Alcubierre metric [1] with $\tanh$-based shaping function $f(r_{s})$ (equation 1) at $v_{s}=0.5$, $R=1$, $\sigma=8$ is the primary test case. Figure 1 shows the NEC comparison between Eulerian and observer-robust evaluation. The Eulerian analysis identifies NEC violations in an annular region at the bubble wall where the energy density is most negative. The observer-robust analysis detects no additional NEC violation points beyond the Eulerian result (Table 3): the violation sets coincide. However, the violation severity differs: the minimum NEC margin across the grid is 0.626-0.626 (robust) versus 0.625-0.625 (Eulerian), a modest difference for the NEC. The WEC tells a more dramatic story (Figure 2): while the violation sets again coincide (0% missed), the minimum WEC margin at $\zeta_{\max}=5$ ($\gamma\approx 74$) is $\sim\!-3450$ for the optimized observer, compared to 0.038-0.038 for the Eulerian frame. This large ratio arises because at NEC-violating points the WEC energy density $T_{ab}u^{a}u^{b}$ scales as $\sim\!\gamma^{2}$ for highly boosted observers (Section 2.4, Interpretation caveat). Nevertheless, this demonstrates that even when the Eulerian frame correctly identifies where violations occur, boosted observers can experience much larger violation magnitudes.
Figure 1: NEC evaluation for the Alcubierre metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1). Left: Eulerian margin (positive = satisfied, negative = violated). Center: Observer-optimized margin. Right: Points where the Eulerian analysis misses violations (red = missed violation). The violation sets coincide (0 missed points), but the robust margin is slightly more negative in the violating region.
Figure 2: WEC evaluation for the Alcubierre metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1). Same layout as Figure 1. The violation sets coincide (0 missed points), but the observer-optimized WEC margin at $\zeta_{\max}=5$ ($\gamma\approx 74$) is much more negative than the Eulerian value; at NEC-violating points this ratio grows with the rapidity cap.
Lentz metric.
The Lentz metric [17] was originally proposed as a “positive-energy” warp drive. However, Celmaster and Rubin [21] showed that the Lentz metric violates the WEC even for Eulerian observers once the metric is properly evaluated. The diamond-pattern shift profile uses $\tanh$-based smooth transitions following the WarpFactory parameterization. Figure 3 shows the NEC comparison. Like the Alcubierre case, the observer-robust analysis finds no additional NEC violations beyond the Eulerian result (Table 3): the six-direction NEC baseline detects the full NEC-violating region at the tested parameters. The NEC violations detected by both analyses are concentrated near the bubble boundary, consistent with the findings of Celmaster and Rubin. Because the Lentz wall spans only $\sim\!0.02$ grid cells at $50^{3}$ (Table 2), the wall is effectively unresolved and any quantitative Lentz results should be regarded as qualitative indicators rather than converged values.
Figure 3: NEC evaluation for the Lentz metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1). NEC violations are concentrated near the bubble boundary. The violation sets coincide between Eulerian and robust analyses (0 missed points).
WarpShell metric.
The WarpShell metric models a constant-velocity physical warp drive following the Bobrick–Martire framework [22] with the Fell–Heisenberg mass parameterization [23], using WarpFactory’s “Bobrick–Martire Modified Time” simplification. It consists of a flat interior ($r<R_{1}$), a curved Schwarzschild-like shell ($R_{1}<r<R_{2}$), and an asymptotically flat exterior ($r>R_{2}$), connected by $C^{2}$-smooth quintic Hermite transition functions (Appendix A). The WarpShell metric uses a shell-like mass distribution (Schwarzschild radius parameter $r_{\mathrm{S}}=5$; smooth_width =0.06=0.06) with an inner approximately flat region, an outer asymptotically flat region (lapse and shift are exactly clamped; the spatial metric retains residual $\tanh$ tails), and a smooth transition layer. The reported WarpShell violations are properties of the regularized numerical implementation, not necessarily of an idealized thin-shell limit; the enormous NEC margin magnitudes ($\mathcal{O}(10^{30})$$\mathcal{O}(10^{35})$) reflect the regularization curvature scale and do not converge with resolution (see Section 8). The spatial metric contains a conformal factor in the shell region: $\gamma_{ij}=\delta_{ij}+(\gamma_{rr}-1)\,\hat{r}_{i}\hat{r}_{j}$ where $\gamma_{rr}$ encodes the shell mass profile. Figure 4 shows the NEC comparison. The WarpShell NEC missed fraction is negligible at $v_{s}=0.5$ (Table 3), and the WEC missed fraction is similarly small (WEC missed $0.1\%$ at $v_{s}=0.5$; DEC missed is negligible at $0.0\%$). Missed WEC violations concentrate at moderate radii ($r\approx 3.5$$4.0$), outside the shell region. The NEC and DEC violations are concentrated at the shell transition boundaries and are likely transition-zone artifacts; the WEC violations at $r\approx 3.5$$4.0$ arise from the residual $\tanh$-tail curvature and represent genuine features of the regularized geometry (see Section 8). For the <1<\!1% non-Type I WarpShell points, the three-term DEC diagnostic (equation 17) is evaluated directly, checking flux causality, future-directedness, and non-negative energy density simultaneously (Section 8).
Figure 4: NEC evaluation for the WarpShell metric (regularized implementation; $50^{3}$ grid, $v_{s}=0.5$, $R_{1}=0.5$, $R_{2}=1.0$; see Table 1). NEC and WEC missed violations are both below 0.1% at $v_{s}=0.5$ (Table 3).
WarpShell uses $C^{2}$-smooth quintic Hermite transitions (see Section 8 for details); the reported violations are properties of this regularized implementation.
Worst-case observer field.
For each grid point, the WEC optimizer returns the worst-case timelike observer parameters $(\zeta^{*},\theta^{*},\phi^{*})$. Figure 5 visualizes this as a boost direction field overlaid on the WEC margin for the Alcubierre metric. The arrows show the spatial direction and magnitude ($|\sinh\zeta^{*}|$) of the Lorentz boost that minimizes the WEC margin. The worst-case observers are predominantly boosted along the direction of bubble propagation, with larger rapidity values near the bubble wall. This is physically intuitive: an observer boosted into the bubble wall “sees” the negative energy density more prominently.
Figure 5: Worst-case WEC observer boost field for the Alcubierre metric ($50^{3}$ grid, $v_{s}=0.5$, $\zeta_{\max}=5$; see Table 1). Arrows show the spatial direction and magnitude ($|\sinh\zeta^{*}|$) of the Lorentz boost that minimizes the WEC margin. The worst-case observers are predominantly boosted along the direction of bubble propagation.
4.2 WarpFactory comparison
As a cross-validation, we compare the Eulerian-only NEC/WEC margins against WarpFactory’s published results [4] for the Alcubierre and Lentz metrics at matched parameters, and then apply the observer-robust analysis.
The Eulerian baseline in warpax evaluates the NEC for six axis-aligned null directions $k^{a}=n^{a}\pm e^{a}_{i}$ for $i\in\{1,2,3\}$, and evaluates WEC/SEC/DEC using the single ADM normal observer $n^{a}$. Our Eulerian results agree with WarpFactory’s published violation regions and qualitative structure; quantitative differences at the level of $\sim\!10^{-3}$$10^{-2}$ relative are expected from the use of autodiff versus fourth-order finite differences for the curvature chain. WarpFactory itself implements observer sampling over $\sim\!1000$ directions [4, 5], going beyond the single-frame Eulerian check. The contribution of warpax is to replace this discrete sampling with continuous observer optimization (Section 2.4), which targets the worst-case observer via gradient descent rather than relying on sampling density.
To quantify the advantage of continuous optimization over discrete sampling, we compare the observer optimizer (8 restarts, up to 256 BFGS iterations each, $\sim\!4000$ objective evaluations per point; see Section 2.4) against dense random observer sampling with rapidity drawn uniformly in $\zeta\in[0,\zeta_{\max}]$ and direction uniform on $S^{2}$ ($\cos\theta\sim U[-1,1]$, $\phi\sim U[0,2\pi)$), at $N_{\text{samp}}=1000,5000,10\,000$ evaluations per point. At $10^{4}$ evaluations, the sampling budget exceeds the optimizer budget ($\sim\!4000$) by a factor of $\sim\!2.5$; the optimizer nevertheless achieves equal or lower minima at the majority of test points and 100% detection, demonstrating that gradient-guided search is more efficient per function evaluation for this problem. This comparison fixes a per-point budget in terms of objective evaluations as a proxy for compute. Each BFGS iteration also requires a reverse-mode autodiff gradient, so an optimizer step is typically several times more expensive than a bare function call. Consequently, an evaluation-count comparison should be interpreted as “efficiency per objective evaluation” rather than a direct wall-clock-time comparison; in wall-clock terms it can otherwise overstate the speed advantage of optimization if gradient cost is not accounted for. We test on 100 randomly selected grid points per metric at $v_{s}=0.5$ and declare a match when the margin difference is $10^{-6}$. For the Rodal metric, the optimizer finds an equal or lower WEC margin at 97% of test points across all sampling densities; for the Alcubierre metric, the rate is 67–69%. In the Alcubierre case, sampling outperforms the optimizer at 31–33% of points, likely because some worst-case observers lie near the rapidity boundary where the $\tanh$-capped parameterization has attenuated gradients, allowing uniform sampling to reach closer to $\zeta_{\max}$. Despite this, the mean optimized margin is lower than the mean sampling margin across all densities.
We define the missed violation fraction for a given energy condition as
$f_{\text{miss}}=\frac{\#\{x\in\text{grid}:m_{\text{Euler}}(x)\geq 0\;\wedge\;m_{\text{rob}}(x)<0\}}{\#\{x\in\text{grid}\}}\,$, (22)
i.e. the fraction of all grid points where the Eulerian margin is non-negative (condition appears satisfied) but the robust ground truth is negative (condition violated). For Type I points (the dominant case; see Table 8), the ground truth $m_{\text{rob}}$ is the algebraic slack (Section 2.2), which is cap-independent; consequently, $f_{\text{miss}}$ is insensitive to $\zeta_{\max}$ in practice. Both $f_{\text{miss}}$ and “% violated” are domain-restricted diagnostics: their values depend on the sampled region and grid density, and are primarily meaningful for comparing the Eulerian and robust analyses within the same domain (not as intrinsic spacetime invariants). Table 3 summarizes $f_{\text{miss}}$ for all metrics at $v_{s}=0.5$. The Rodal metric shows substantial missed violations across NEC, WEC, SEC, and DEC. SEC misses are notably widespread: Van Den Broeck has $1.2\%$ SEC missed despite nonzero NEC/WEC/DEC misses, demonstrating that SEC violations are genuinely independent of the other conditions. This independence arises because SEC involves the trace $\rho+\sum_{i}p_{i}$, which can be negative even when each individual $\rho+p_{i}\geq 0$ (NEC satisfied). The violation severity can differ substantially between Eulerian and optimized observers even when the violation sets coincide. Table values are percentages over the full $50^{3}$ volume rounded to $0.1\%$; slice plots may show a few points even when the rounded volume fraction is $0.0\%$. In unrounded values, WarpShell shows NEC total violated at $0.259\%$ and WEC total violated at $0.270\%$. A slight WEC >> NEC difference is expected because the WEC is strictly stronger than the NEC: WEC satisfaction implies NEC satisfaction, but WEC can be violated even when NEC holds. At this sub-percent level, exact counts can also be influenced by finite-precision effects near the violation boundary and by the fact that our non-Type I diagnostics are $\zeta_{\max}$-capped. At the 0.1% reporting precision of Tables 3 and 4, both values round to the same figure; the underlying sub-percent difference is visible only at higher precision. Table 4 shows the velocity dependence for the two metrics with the largest discrepancies.
Table 3: Energy condition analysis at $v_{s}=0.5$ on a $50^{3}$ grid ($\zeta_{\max}=5$ throughout). Total: percentage of grid points violating each condition (hybrid margin: algebraic slacks at Type I points, $\zeta_{\max}$-capped extrema at non-Type I). Missed: percentage of grid points where the Eulerian analysis fails to detect a violation found by the hybrid check. The Eulerian violated fraction equals the total violated fraction minus the missed fraction. A missed fraction of $0.0\%$ does not imply zero violations; it means all violations are detected by the Eulerian frame. Bold entries highlight missed fractions $\geq 1\%$. Lentz and WarpShell are under-resolved ($\ll\!1$ and $\sim\!0.3$ cells across the wall, respectively; Table 2); their total and missed fractions should be interpreted as lower bounds.
Total violated (%) Missed by Eulerian (%)
Metric NEC WEC SEC DEC NEC WEC SEC DEC
Schwarzschild 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Alcubierre 2.1 2.1 2.1 2.1 0.0 0.0 0.3 0.0
Lentz 0.1 0.1 0.1 0.1 0.0 0.0 0.1 0.0
Van Den Broeck 6.5 6.5 6.6 6.9 0.1 0.4 1.2 0.3
Natário 2.5 2.5 2.5 2.5 0.0 0.0 0.0 0.0
Rodal 87.0 87.0 87.0 99.9 1.6 15.6 28.0 28.5
WarpShell 0.3 0.3 0.3 0.3 0.0 0.1 0.1 0.0
Table 4: Velocity dependence of missed violations for Rodal and WarpShell metrics on a $50^{3}$ grid ($\zeta_{\max}=5$). Rodal misses are persistent across velocities; WarpShell missed fractions are uniformly below $0.1\%$ for all conditions and velocities.
Metric $v_{s}$ NEC (%) WEC (%) SEC (%) DEC (%)
Rodal 0.10 1.3 14.9 27.3 28.5
0.50 1.6 15.6 28.0 28.5
0.90 1.6 15.6 28.0 28.5
0.99 1.6 15.6 28.0 28.5
WarpShell 0.10 0.0 0.1 0.1 0.0
0.50 0.0 0.1 0.1 0.0
0.90 0.0 0.1 0.1 0.0
0.99 0.0 0.0 0.1 0.0
4.3 Velocity dependence
We study the dependence of energy condition violations on the bubble velocity $v_{s}$ across all six metrics, sweeping over $v_{s}\in\{0.1,0.5,0.9,0.99\}$.
For the Alcubierre metric (Figure 7a), the Eulerian frame detects all NEC and WEC violation points at every velocity tested (missed $0.0\%$). The robust NEC violation fraction grows from 1.73% at $v_{s}=0.1$ to 2.21% at $v_{s}=0.99$, and the robust WEC violation fraction tracks NEC identically; the violation regions coincide with the Eulerian analysis, but the observer-optimized margins are consistently more negative, indicating greater violation severity for boosted observers.
For the Rodal metric, the discrepancy is far more dramatic. The missed DEC violation fraction is approximately 28.5% at all velocities, indicating that the Eulerian DEC analysis fails to detect violations at roughly 28.5% of all grid points. Missed WEC violations reach  ${\sim}\,15\%$ across the velocity sweep, growing from 14.9% at $v_{s}=0.1$ to 15.6% at $v_{s}=0.99$. Even the NEC missed fraction is non-zero for Rodal, ranging from 1.26% to 1.61%. The algebraic DEC slacks at missed points are shallow: the observer-optimized minimum is $2.0\times 10^{-6}$ (Section 8), and deterministic Fibonacci sampling saturates at $\sim\!93$% detection even at $10^{4}$ directions per point (Figure 13), confirming that the undetected 28.5% comprises violations near the satisfaction boundary rather than deeply negative points.
Figure 6 shows the spatial distribution of missed violations at each velocity. The central result is that the discrepancy between Eulerian and observer-robust analyses is systematic and persistent across the full velocity range, with the strongest effect in the DEC and WEC conditions rather than the NEC.
Figure 6: Energy condition violations missed by Eulerian analysis ($50^{3}$ grid, $\zeta_{\max}=5$) across four bubble velocities $v_{s}\in\{0.1,0.5,0.9,0.99\}$. Red regions indicate points where a condition appears satisfied for the Eulerian observer but is violated for the worst-case observer. The missed violation fraction is most pronounced for the DEC and WEC conditions.
Figure 7: (a) Minimum NEC margin as a function of bubble velocity for the Alcubierre metric ($50^{3}$ grid; see Table 1). The Eulerian margin (solid, circles) and observer-robust margin (dashed, squares) track closely, confirming that the Eulerian frame captures the NEC violation extent for this geometry. Both margins become more negative with increasing $v_{s}$. (b) Richardson extrapolation resolution-stability study for the Alcubierre metric NEC minimum margin ($25^{3}$$50^{3}$$100^{3}$ grids). Data points (circles) and reference $p=2$ slope (dashed); the three-point data is non-monotone, so the line indicates the assumed order, not a measured convergence rate. Error is measured relative to the Richardson-extrapolated value.
5 Results: Geodesics
5.1 Tidal forces
We integrate a timelike geodesic through the Alcubierre bubble at $v_{s}=0.5$, starting at $x=5R$ with initial spatial 4-velocity components $u^{x}=-0.3$, $u^{y}=u^{z}=0$ (a small offset $y=0.1$ avoids the symmetry axis). The temporal component $u^{0}$ is determined by solving the normalization constraint $g_{\mu\nu}\,u^{\mu}u^{\nu}=-1$ for the given spatial components; integration uses proper time $\tau$ as the affine parameter. At each saved point along the trajectory, the tidal tensor $K^{\mu}{}_{\rho}=R^{\mu}{}_{\nu\rho\sigma}\,v^{\nu}v^{\sigma}$ is evaluated via autodiff. This $4\times 4$ mixed tensor has one zero eigenvalue (along $v^{\mu}$); the three non-zero eigenvalues are extracted and correspond to the eigenvalues of the spatial tidal matrix $\mathcal{E}_{ij}=R_{\hat{0}i\hat{0}j}$ in an orthonormal comoving frame (see Section 2.5).
Figure 8 shows the three independent tidal eigenvalues as a function of proper time. As the observer crosses the bubble wall, two eigenvalues spike sharply, indicating strong tidal compression in the transverse directions, while the third shows a corresponding stretching along the radial direction. Inside the bubble, the tidal forces are numerically negligible (consistent with a nearly flat interior for smooth shaping functions). The peak tidal force increases with both bubble velocity and wall steepness $\sigma$ (larger $\sigma$ corresponds to a thinner, steeper wall), consistent with the analytic expectation that thinner bubble walls produce more extreme tidal effects.
Figure 8: Tidal eigenvalue evolution along a radial timelike geodesic through the Alcubierre bubble ($v_{s}=0.5$, $R=1$, $\sigma=8$; ODE integration via Tsit5; see Table 1). The three eigenvalues of the tidal tensor $K^{\mu}{}_{\rho}$ are plotted as a function of proper time. Sharp tidal forces at the bubble wall indicate significant stretching and compression.
5.2 Blueshift analysis
Photon blueshift at the bubble wall is computed by integrating a null geodesic from the far field through the bubble and evaluating the frequency ratio (equation 21) with respect to coordinate-stationary observers ($u^{a}=(1/\sqrt{|g_{00}|},0,0,0)$) at each point. The initial null 4-momentum is constructed by choosing a spatial propagation direction $n^{i}$ (here $n^{i}=(-1,0,0)$, toward the bubble) and solving the null constraint $g_{\mu\nu}\,k^{\mu}k^{\nu}=0$ for $k^{0}$ via the quadratic formula, selecting the future-directed root. Since null geodesics have a free affine rescaling $k^{\mu}\to\lambda k^{\mu}$, no further normalization is imposed; the frequency ratio in equation (21) is independent of this choice. Figure 9 shows the blueshift profile along the photon trajectory. A photon approaching the leading edge of the bubble experiences a blueshift that grows monotonically with velocity: the peak frequency ratio is $1.005$ at $v_{s}=0.1$, rising to $1.155$ at $v_{s}=0.5$, $2.294$ at $v_{s}=0.9$, and $7.089$ at $v_{s}=0.99$. These values match the Lorentz factor $\gamma(v_{s})=1/\sqrt{1-v_{s}^{2}}$ to four significant figures ($\gamma=1.005$, $1.155$, $2.294$, $7.089$ for the respective velocities). This is expected: in the Alcubierre bubble interior where $f\approx 1$, the coordinate-stationary observer has proper time $\mathrm{d}\tau=\sqrt{1-v_{s}^{2}f^{2}}\,\mathrm{d}t$, giving a characteristic frequency ratio of $\gamma(v_{s})$. More precisely, the constant-time Alcubierre metric with lapse $\alpha=1$ admits the helical Killing vector $K=\partial_{t}+v_{s}\partial_{x}$; we observe four-significant-figure agreement with $\gamma(v_{s})$ at all tested velocities, consistent with a Killing-vector argument in comoving coordinates (a full derivation is deferred to future work). This agreement provides strong numerical validation of the geodesic integrator. Photons at the trailing edge are redshifted. This asymmetry reflects the Doppler-like effect of the moving spacetime distortion on photon propagation.
Figure 9: Photon frequency ratio along a null geodesic through the Alcubierre bubble ($v_{s}=0.5$, $R=1$, $\sigma=8$; ODE integration via Tsit5; see Table 1). The blueshift at the leading edge and redshift at the trailing edge reflect the asymmetric spacetime distortion.
6 Results: Kinematic scalars
The kinematic decomposition of the Eulerian congruence provides complementary information about the spacetime geometry. The expansion $\theta$, shear scalar $\sigma^{2}=\sigma_{ab}\,\sigma^{ab}$, and vorticity scalar $\omega^{2}=\omega_{ab}\,\omega^{ab}$ are computed from the extrinsic curvature $K_{ij}$ of the spatial hypersurfaces [3]:
$\theta=-K^{i}{}_{i}\,,\quad\sigma_{ij}=-K_{ij}-\frac{1}{3}\theta\,\gamma_{ij}\,,\quad\omega_{ij}=0\;\text{(hypersurface-orthogonal)}\,.$ (23)
The extrinsic curvature is computed from the Christoffel symbols via $K_{ij}=-\alpha\,\Gamma^{0}{}_{ij}$, where $\alpha$ is the ADM lapse. For the Eulerian congruence (which is hypersurface-orthogonal by construction in the ADM decomposition), the vorticity vanishes identically.
Figure 10 shows the kinematic scalars for the Alcubierre metric at $v_{s}=0.5$. The expansion $\theta$ has the characteristic bipolar structure: positive expansion ahead of the bubble (space expanding) and negative expansion behind (space contracting), with $\theta=0$ inside the flat region. The shear is concentrated at the bubble wall and peaks where the expansion gradient is largest. These kinematic features are closely connected to the exotic matter requirements. For a geodesic, irrotational timelike congruence, the Raychaudhuri equation $\dot{\theta}=-\tfrac{1}{3}\theta^{2}-\sigma_{ab}\sigma^{ab}-R_{ab}\,u^{a}u^{b}$ links the expansion evolution to $R_{ab}\,u^{a}u^{b}$, and hence (via the Einstein equations) to the SEC. The Eulerian congruence is in general accelerated (with acceleration $a^{b}=\nabla_{b}\ln\alpha$), so the connection is qualitative rather than a strict implication; a similar link to the NEC holds for null geodesic congruences.
Figure 10: Kinematic scalars for the Alcubierre metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1): expansion $\theta$ (left), shear $\sigma^{2}$ (center), and vorticity $\omega^{2}$ (right). The bipolar expansion structure is the hallmark of the warp drive geometry.
Figure 11 shows the same scalars for the Lentz metric. Despite the different shift vector construction, the expansion structure is qualitatively similar to Alcubierre’s, with the key difference that the Lentz metric was designed to distribute the expansion more smoothly, reducing the peak negative expansion. However, this smoother distribution does not eliminate NEC violations, which are present even in the Eulerian frame (Section 4.1).
Figure 11: Kinematic scalars for the Lentz metric ($50^{3}$ grid, $v_{s}=0.5$; see Table 1). While the expansion is more smoothly distributed than the Alcubierre case, the qualitative bipolar structure persists.
7 Resolution stability
We validate the numerical accuracy of the curvature chain and energy condition evaluation using Richardson extrapolation at three grid resolutions: $25^{3}$, $50^{3}$, and $100^{3}$ (Appendix C collects additional per-metric wall-resolution data). The Richardson extrapolation assumes $p$th-order convergence of the error with grid spacing $h=1/N$, where $N$ is the number of grid points along each axis:
$Q(h)=Q_{\text{exact}}+C\,h^{p}+\mathcal{O}(h^{p+1})\,$, (24)
where $Q$ is a scalar quantity of interest (e.g., the minimum NEC margin over the grid). Using $h=1/N$ (rather than L/(N1)L/(N{-}1)) as the spacing proxy gives an exact refinement ratio $r=2$ for $N=25,50,100$; the $\sim\!2\%$ difference from the node-to-node spacing L/(N1)L/(N{-}1) is smaller than the convergence error at these resolutions. From evaluations at three resolutions we extract the convergence order $p$ and the extrapolated value $Q_{\text{exact}}$.
For the Alcubierre metric, the three-point sequences for both the minimum NEC margin (0.631,0.626,0.629-0.631,\,-0.626,\,-0.629) and the integrated violation volume (1.811, 1.905, 1.7441.811,\,1.905,\,1.744) are non-monotone, so the standard Richardson ratio test does not yield a meaningful observed convergence order. Assuming second-order convergence ($p=2$), the Richardson-extrapolated values differ from the $100^{3}$ results by $\sim\!0.1$% (minimum margin) and $\sim\!3$% (integrated violation), confirming resolution stability. We emphasize that these are resolution-stability checks for discrete summaries (a grid-sampled minimum and a thresholded volume), not formal asymptotic error guarantees; the goal is to verify that the autodiff curvature computation introduces no additional discretization beyond spatial sampling. The unnormalized $L^{2}$-norm $\sqrt{\sum_{i}f_{i}^{2}}$ of the NEC violation field does not converge (observed $p\approx-1.3$) because this discrete sum is an extensive quantity that grows with the number of grid points at fixed domain extent; a Riemann-sum approximation $\sqrt{\sum_{i}f_{i}^{2}\,h^{3}}$ (with cell volume $h^{3}$) would be needed for convergence to the integral $\|f\|_{L^{2}}$. The global minimum is not a smooth functional of the discrete field; the non-monotone three-point sequence is consistent with second-order behavior for the $50^{3}\to 100^{3}$ pair, but the $25^{3}$ point lies outside the asymptotic regime for this statistic. This convergence study validates grid sampling of scalar summaries (minimum margin, integrated violation volume), not the curvature operator itself; the latter is exact (up to floating-point roundoff) via autodiff at each grid point for the implemented metric. Figure 7b shows the convergence log-log plot and Table 5 summarizes the numerical results.
Table 5: Richardson extrapolation data for the Alcubierre NEC on $25^{3}$$50^{3}$$100^{3}$ grids. The minimum margin and integrated violation are resolution-stable (assumed $p=2$; three-point data is non-monotone, indicating that $25^{3}$ lies outside the asymptotic regime for these statistics); the $L^{2}$ norm does not converge (extensive quantity).
Quantity N=25N\!=\!25 N=50N\!=\!50 N=100N\!=\!100 Extrap. $p$ Error est.
Min margin NEC 0.631-0.631 0.626-0.626 0.629-0.629 0.631-0.631 $(2)^{\dagger}$ $1.1\times 10^{-3}$
Integrated viol. $1.811$ $1.905$ $1.744$ $1.691$ $(2)^{\dagger}$ $5.3\times 10^{-2}$
$L^{2}$ viol. norm $3.00$ $8.77$ $23.2$ 1.3-1.3
The assumed second-order extrapolation provides a basic resolution-stability check: the $50^{3}$ value used for the main results is within $\sim$1% of the extrapolated value for the minimum NEC margin. The integrated violation volume has a larger error at $50^{3}$ (closer at $100^{3}$), consistent with the need for finer grids to resolve sharp violation boundaries. Because the reported quantities involve a discrete minimum and a thresholded volume, and the three-point data is non-monotone, the “order” should be interpreted as an assumption for extrapolation rather than a measured convergence rate.
Assumed; three-point data is non-monotone (standard Richardson ratio test inapplicable).
Resolution stability for Rodal and WarpShell.
Table 2 characterizes the wall resolution for each metric at $50^{3}$. Rodal is the only metric that comfortably exceeds the 4-cell resolution threshold ($\sim\!6$ cells); Alcubierre, VdB, and Natário are marginally resolved ($\sim\!1.4$ cells), Lentz is severely under-resolved ($\sim\!0.02$ cells), and WarpShell concentrates its thin-shell transition within $\lesssim\!1$ cell. Autodiff computes exact curvature at each grid point regardless of wall resolution. To verify that the headline missed-violation fractions are not resolution artifacts, we run the same three-resolution sweep ($25^{3}$$50^{3}$$100^{3}$) for the Rodal ($R=100$, $\sigma=0.03$) and WarpShell ($R_{1}=0.5$, $R_{2}=1.0$) metrics. At $100^{3}$ we use Eulerian-only evaluation due to the prohibitive cost of full observer optimization at $10^{6}$ points. The minimum NEC margin converges at second order for Rodal: extrapolation gives $2.58\times 10^{-6}$ (503 value: $2.62\times 10^{-6}$, within 1.5%). For WarpShell, the Eulerian minimum NEC margin increases in absolute magnitude with resolution ($2.5\times 10^{25}$ at $25^{3}$, $2.3\times 10^{33}$ at $50^{3}$, $2.1\times 10^{34}$ at $100^{3}$), reflecting the shell boundary extremes that are better resolved at finer grids; Richardson extrapolation yields a negative convergence order ($p\approx-3.0$), so we report per-resolution values rather than an extrapolated limit. The Rodal integrated violation does not converge monotonically, consistent with the broad Rodal wall ($\sigma=0.03$, 10–90% transition width $73.24$ coordinate units spanning $\sim\!6.0$ grid cells; Table 2), which is marginally resolved at the $50^{3}$ grid but whose small gradients require higher resolution to stabilize extensive (volume-integrated) quantities.
Rodal missed-fraction stability.
To verify that the Rodal missed-violation percentages reported in Table 3 are not artifacts of the $50^{3}$ grid, we evaluate the full Eulerian-vs-robust comparison at three resolutions ($25^{3}$$50^{3}$$100^{3}$) using $N_{\text{starts}}=8$. Since Rodal is 100% Type I at all three resolutions (24 non-Type I points at $25^{3}$ due to coarse-grid effects), the algebraic slacks provide exact truth. Table 6 shows that the missed fractions are stable: NEC miss decreases monotonically from 1.8% to 1.5%, WEC miss is constant at $\sim\!15.6\%$, SEC miss from 28.4% to 27.8%, and DEC miss is stable at $\sim\!28.5\%$. The total violation fractions are likewise stable ($\sim\!87$% for NEC/WEC/SEC, $\sim\!99.9$% for DEC). These results confirm that the headline missed-violation fractions are resolution-independent at the reported grid spacing.
Table 6: Rodal missed-fraction stability across grid resolutions ($v_{s}=0.5$, $R=100$, $\sigma=0.03$, $N_{\text{starts}}=8$). All points are Type I at $N\geq 50$; truth is the algebraic slack. Missed fractions vary by <2<2 percentage points across a $64\times$ increase in grid volume.
$f_{\text{miss}}$ (%) Total violated (%)
NN NEC WEC SEC DEC NEC WEC SEC DEC
25 1.8 15.6 28.4 28.2 87.2 87.2 87.2 99.8
50 1.6 15.6 28.0 28.5 87.0 87.0 87.0 100.0
100 1.5 15.6 27.8 28.7 86.9 86.9 86.9 99.9
Coordinate-invariance check.
As a sanity check on the curvature computation, we verify the traveling-wave property of the Alcubierre metric. Two bubble configurations with $x_{s}=0$ and $x_{s}=1$ (corresponding to the bubble center at $t=0$ and $t=1/v_{s}$) are evaluated on a $25^{3}$ grid spanning $5,5]^{3}$. The center of mass of the NEC violation pattern shifts by $1.002$ in $x$ (expected $1.000$), with transverse shifts <0.01<0.01. Violation counts differ by <3<3% (a discrete-grid sampling effect), and the minimum NEC margins agree to machine precision. This confirms that the energy condition diagnostics respect the spatial translation symmetry of the Alcubierre metric.
7.1 Curvature identity validation
As an independent check of the autodiff curvature chain, we verify that the computed Riemann tensor satisfies fundamental geometric identities at representative points for the Schwarzschild and Alcubierre metrics. The three algebraic Riemann symmetries (antisymmetry in the first and second index pairs, $R_{abcd}=-R_{bacd}$ and $R_{abcd}=-R_{abdc}$, and pair interchange symmetry, $R_{abcd}=R_{cdab}$) hold with maximum residual $10^{-12}$ across all test points. The first Bianchi identity ($R_{abcd}+R_{acdb}+R_{adbc}=0$) is satisfied to the same precision. The contracted Bianchi identity ($\nabla_{a}G^{ab}=0$) is verified to a residual $10^{-6}$ (absolute, normalized by $\|G^{ab}\|_{\max}$ this corresponds to a relative residual $\lesssim 10^{-6}$) via autodiff evaluation of the covariant divergence; the larger residual ($10^{-6}$ vs. $10^{-12}$ for the algebraic identities) is expected because this identity involves third derivatives of the metric, where floating-point cancellation is more severe. All computations use 64-bit floating point (jax.config.update("jax_enable_x64", True)). Since $T_{ab}=G_{ab}/(8\pi)$ and $G_{ab}$ inherits the symmetry of $R_{ab}$ and $g_{ab}$, the stress-energy tensor is symmetric by construction; spot-checked residuals $|T_{ab}-T_{ba}|_{\max}<10^{-14}$ confirm this numerically. These results confirm that the forward-mode autodiff chain correctly preserves the geometric structure of the curvature tensor.
7.2 Optimization stability
To verify that the multi-start observer optimization converges reliably, we perform an ablation study over the number of random restarts $N_{\text{starts}}\in\{1,2,4,8,16\}$ for three representative metrics at $v_{s}=0.5$: Alcubierre (zero missed violations, severity only), Rodal (largest missed fraction), and WarpShell (negligible missed violations, confirming regularized-metric stability). For each metric, the curvature grid is computed once and only the observer optimization is re-run at each $N_{\text{starts}}$ value. The restart sets at different $N_{\text{starts}}$ use independent random draws (not nested subsets), so the observed stability reflects robustness to both initialization count and seed variation. All multi-start initializations use JAX’s splittable PRNG with a fixed base key, producing deterministic results on a given hardware backend. Results are not bitwise reproducible across CPU and GPU due to floating-point non-associativity in parallel reductions.
Table 7 shows the results. For all three metrics, the missed violation percentages stabilize by $N_{\text{starts}}=2$. For Alcubierre, $f_{\text{miss}}=0$ at all counts; the minimum hybrid margin jumps from 0.14-0.14 at $N_{\text{starts}}=1$ to 3449-3449 at $N_{\text{starts}}=2$ as the optimizer discovers deeper violations at non-Type I transition-zone points, and remains converged thereafter. For Rodal (100% Type I), the algebraic slacks are exact and optimizer-independent; the missed WEC fraction is constant at 15.60% across the full sweep. The WarpShell metric has <1<\!1% non-Type I points at the shell boundary, and the minimum WEC margin converges to six significant figures by $N_{\text{starts}}=4$. These results confirm that $N_{\text{starts}}=8$, used throughout this paper, provides converged results with a safety margin.
Table 7: $N_{\text{starts}}$ ablation at $v_{s}=0.5$ on a $50^{3}$ grid ($\zeta_{\max}=5$). Rodal is 100% Type I; Alcubierre $\sim\!98$%; WarpShell $\sim\!99.6$%. The missed WEC fraction $f_{\text{miss}}$ is constant for Rodal (100% Type I) and stabilizes from $N_{\text{starts}}\geq 2$ for Alcubierre and WarpShell, confirming that violation detection is robust to optimizer configuration once a minimal multi-start budget is met. Minimum-margin estimates are likewise stable from $N_{\text{starts}}\geq 2$ for Alcubierre and from $N_{\text{starts}}\geq 4$ for WarpShell, consistent with the small (<1<\!1%) non-Type I subset driving the remaining multi-start sensitivity.
$N_{\text{starts}}$
Metric 1 2 4 8 16
Min hybrid WEC margin
Alcubierre 0.14-0.14 3449-3449 3449-3449 3449-3449 3449-3449
Rodal $2.6\!{\times}\!10^{-6}$ $2.6\!{\times}\!10^{-6}$ $2.6\!{\times}\!10^{-6}$ $2.6\!{\times}\!10^{-6}$ $2.6\!{\times}\!10^{-6}$
WarpShell $4.8\!{\times}\!10^{35}$ $3.4\!{\times}\!10^{35}$ $3.7\!{\times}\!10^{35}$ $3.7\!{\times}\!10^{35}$ $3.5\!{\times}\!10^{35}$
Missed WEC (%)
Alcubierre 0.00 0.00 0.00 0.00 0.00
Rodal 15.60 15.60 15.60 15.60 15.60
WarpShell 0.06 0.04 0.05 0.05 0.05
Hybrid margin: algebraic slack at Type I points; capped extremum at non-Type I points.
Optimizer convergence.
To assess observer optimizer reliability, we examine per-point convergence diagnostics for the WarpShell metric at $v_{s}=0.5$ (the only metric for which per-point convergence data was recorded in the current analysis run). The NEC optimizer converges at 99.8% of grid points with a median of 2 iterations; 0.23% of points reach the maximum iteration limit of 256. The WEC and SEC optimizers converge at 99.7% with 0.31% hitting the iteration limit, while the DEC converges at 99.6% with 0.39% at the limit. Across all four conditions, the median iteration count is 2 and the mean is below 3, indicating that the optimization landscape is well-conditioned at the vast majority of grid points. The slightly lower DEC convergence rate is consistent with the DEC objective combining flux causality, future-directedness, and energy density, a more complex landscape than the single-contraction NEC or WEC objectives. These results confirm that the multi-start observer optimization procedure converges reliably, with fewer than 0.4% of points requiring the full iteration budget.
7.3 Classification breakdown and conditional miss rate
Table 8 reports the Hawking–Ellis type breakdown for all seven metrics at $v_{s}=0.5$ (Schwarzschild at $v_{s}=0$). We employ a two-tier imaginary tolerance (Section 8): (i) absolute, $|\mathrm{Im}\,\lambda_{i}|<\tau\cdot s$ with $s=\max(|\mathrm{Re}\,\lambda|,1)$, and (ii) relative, $|\mathrm{Im}\,\lambda_{i}|<3\times 10^{-3}\cdot\max|\mathrm{Re}\,\lambda|$ (unclamped), which catches split degenerate pairs at large $\|T\|$. Rodal, Schwarzschild, and Lentz are $\geq\!99.9$% Type I. Alcubierre, Natário, and VdB are $\sim\!97$9898% Type I, with the remaining points classified as Type IV at the transition zone between the warp bubble and flat space. WarpShell achieves $\sim\!99.6$% Type I, with $\sim\!0.3$% Type II at the shell boundary and $\sim\!0.1$% Type IV. WarpShell serves primarily as a numerical regularity stress test: its prescribed $C^{2}$ transition functions generate extreme curvature scales that exercise the classifier, optimizer, and convergence diagnostics in a worst-case regime; the resulting EC margins should be understood as properties of the regularized implementation, not as predictions for an idealized thin-shell spacetime. The residual non-Type I WarpShell points are concentrated in the shell transition region where $\|T^{a}{}_{b}\|$ reaches $\mathcal{O}(10^{25})$ (at $50^{3}$ resolution; the eigenvalue scale grows with resolution as the grid resolves the thin shell). A dedicated tetrad-normalization diagnostic confirms that the observer-orthonormal tetrad is well-normalized ($\max|g(k,k)|<10^{-6}$ and $\max|\eta_{IJ}-g(e_{I},e_{J})|<10^{-6}$ at >99.8>99.8% of grid points), and that NEC contractions $T_{ab}k^{a}k^{b}$ are the same order of magnitude as the eigenvalues of $T^{a}{}_{b}$ (ratio $\sim 0.4$). At finer grids the eigenvalue scale itself increases (reaching $\sim\!10^{33}$ at $100^{3}$; Section 7), and observer boosting at $\zeta_{\max}=5$ ($\gamma\approx 74$) further amplifies contractions by $\cosh^{2}\!\zeta\approx 5500$, bridging the gap between the $50^{3}$ eigenvalue scale and the reported $\mathcal{O}(10^{30})$$\mathcal{O}(10^{35})$ margins. These therefore reflect genuine curvature amplitudes at thin-shell grid points, not a normalization artefact. At these points, eigenvalue pairs of the non-symmetric mixed tensor split as $\lambda\pm\varepsilon i$ with $|\varepsilon|/|\lambda|\sim 10^{-3}$$10^{-2}$. The two-tier tolerance (Section 8) recovers the majority of such points as effectively real; the residual 0.10.1% Type IV fraction is comparable to that of the other warp metrics. The classifier is conservative by design: a point is labeled Type I only when eigenvalues pass either tier of the imaginary-part check. Any ambiguous point defaults to the optimizer pathway. Misclassifying Type I as non-Type I incurs only a performance penalty (unnecessary optimization); the reverse (labeling non-Type I as Type I) would require complex eigenvalues to simultaneously satisfy both the absolute and relative tolerance, which the tolerance sensitivity analysis (Section 8) confirms is extremely unlikely at the tested thresholds. At non-Type I points the capped extrema are used; at Type I points the algebraic slacks dominate.
Table 8: Hawking–Ellis type breakdown at $v_{s}=0.5$ ($50^{3}$ grid; Schwarzschild at $v_{s}=0$). Metrics with steep shape functions (Natário, VdB) show slightly higher non-Type I fractions due to numerically split degenerate eigenvalues at transition zones.
Metric % Type I % Type II % Type IV max $|\mathrm{Im}\,\lambda|/s$
Rodal 100.0 0.0 0.0 0
Schwarzschild 100.0 0.0 0.0 $7.3\times 10^{-19}$
Lentz 99.9 0.0 0.1 $2.8\times 10^{-1}$
Alcubierre 98.0 0.0 2.0 $2.7\times 10^{-1}$
Natário 97.5 0.0 2.5 $1.7\times 10^{0}$
VdB 98.4 0.0 1.6 $2.2\times 10^{-1}$
WarpShell 99.6 0.3 0.1 $1.4\times 10^{34}$
The WarpShell $\max|\mathrm{Im}\,\lambda|/s$ is dominated by rare near-zero-$\mathrm{Re}\,\lambda$ outliers where $s=\max(|\mathrm{Re}\,\lambda_{j}|,1)=1$; typical split ratios are $10^{-3}$$10^{-2}$, consistent with the two-tier tolerance analysis below.
We additionally define the conditional miss rate $f_{\text{miss}|\text{viol}}$ as the fraction of robust-violated points missed by the Eulerian analysis:
$f_{\text{miss}|\text{viol}}=\frac{\#\{\text{Eulerian-satisfied}\cap\text{robust-violated}\}}{\#\{\text{robust-violated}\}}\times 100\%.$ (25)
This quantity measures how “leaky” the Eulerian frame is at points where violations actually exist, complementing the unconditional $f_{\text{miss}}$ which is diluted by the (potentially large) number of satisfied points.
8 Discussion
The central result of this work is that single-frame energy condition evaluation can systematically underestimate the extent of energy condition violations in warp drive spacetimes, though the severity depends strongly on the metric geometry. For the Rodal metric, the Eulerian DEC analysis fails to detect violations at over 28% of all grid points, and the WEC figure exceeds 15%. For WarpShell (regularized implementation), missed WEC violations are below $0.1\%$ of the grid. For the Alcubierre and Natário metrics, we find no additional NEC or WEC violating grid points beyond the baseline Eulerian analysis at the tested parameters (missed =0=0 at the sampled resolution). Lentz shows zero missed NEC/WEC points (though nonzero SEC misses exist) and is severely under-resolved ($\ll\!1$ cell; Table 2); this result should be treated as a lower bound. Van Den Broeck exhibits small but nonzero missed fractions (NEC $0.1\%$, WEC $0.4\%$, DEC $0.3\%$ at $v_{s}=0.5$), and non-trivial SEC misses persist, most notably $1.2\%$ for Van Den Broeck (Table 3). Even for these metrics the observer-optimized violation severity can be much larger (Section 2.4, Interpretation caveat): for the Alcubierre WEC at $\zeta_{\max}=5$, the capped minimum is $\sim\!-3400$ compared to 0.038-0.038 for the Eulerian frame. These severity ratios are $\zeta_{\max}$-dependent and grow without bound; single-frame analysis can underestimate the magnitude of violations even when it correctly identifies their spatial extent.
Table 9 quantifies this dependence for the Alcubierre metric at $v_{s}=0.5$ on a $25^{3}$ grid. The algebraic slack (“Alg. min”), computed as the minimum eigenvalue-based WEC slack over Type I points only, is $\zeta_{\max}$-independent at $\sim\!{-5}\times 10^{-11}$ (essentially zero). The Eulerian margins and the NEC robust margins are likewise cap-independent. The capped observer extremum (“Capped min”), however, deepens from $\sim\!-1.1$ at $\zeta_{\max}=1$ to $\sim\!-1.9\times 10^{5}$ at $\zeta_{\max}=7$, scaling approximately as $e^{2\zeta_{\max}}$, consistent with the expected $\gamma^{2}$ scaling at NEC-violating Type I points when boosting toward the null direction. The capped minimum tracks the same spatial point (an NEC-violating transition-zone point) across all caps; the deepening reflects genuine $\gamma^{2}$ boosting, not near-vacuum numerical noise. The missed-violation fraction $f_{\text{miss}}$ remains exactly $0.03\%$ across all caps, confirming that violation detection is $\zeta_{\max}$-independent.
Table 9: Rapidity-cap sensitivity for the Alcubierre metric at $v_{s}=0.5$ on a $25^{3}$ grid ($n_{\text{starts}}=8$). “Alg. min” is the minimum algebraic WEC slack over Type I points only, and is $\zeta_{\max}$-independent at $\sim\!{-5}\times 10^{-11}$. “Capped min” is the minimum capped extremum across all points (including non-Type I), which deepens with the cap as $\sim\!e^{2\zeta_{\max}}$, reflecting $\gamma^{2}$ boosting at NEC-violating points. Eulerian margins, NEC robust margins, and $f_{\text{miss}}$ are likewise cap-independent.
$\zeta_{\max}$ WEC NEC
Eul. min Alg. min Capped min $f_{\text{miss}}$ Eul. min Rob. min $f_{\text{miss}}$
1 $3.1\times 10^{-2}$ $\sim\!{-5}\times 10^{-11}$ $1.1\times 10^{0\phantom{0}}$ 0.03% $6.3\times 10^{-1}$ $6.3\times 10^{-1}$ 0.00%
3 $3.1\times 10^{-2}$ $\sim\!{-5}\times 10^{-11}$ $6.4\times 10^{1\phantom{0}}$ 0.03% $6.3\times 10^{-1}$ $6.3\times 10^{-1}$ 0.00%
5 $3.1\times 10^{-2}$ $\sim\!{-5}\times 10^{-11}$ $3.5\times 10^{3\phantom{0}}$ 0.03% $6.3\times 10^{-1}$ $6.3\times 10^{-1}$ 0.00%
7 $3.1\times 10^{-2}$ $\sim\!{-5}\times 10^{-11}$ $1.9\times 10^{5\phantom{0}}$ 0.03% $6.3\times 10^{-1}$ $6.3\times 10^{-1}$ 0.00%
Physical interpretation.
For the Alcubierre metric at $v_{s}=0.5$, $R=1$, $\sigma=8$ (Figure 5), we observe empirically that the worst-case WEC observers are predominantly Lorentz-boosted along the direction of bubble propagation. This is physically plausible: boosted observers see a different effective stress-energy distribution, and the Eulerian frame, adapted to the spatial hypersurface, is not aligned with the worst-case direction. For the Rodal DEC, the alignment angle between the optimizer’s boost direction and the DEC-worst eigenvector spatial direction has a median of $82^{\circ}$ across all DEC-violating Type I points at four velocities (Figure 12), with only 0.90.9% of points having alignment below $10^{\circ}$. This confirms that the worst-case DEC observers are overwhelmingly non-Eulerian and that single-frame analysis misses violations precisely because it fails to probe these nearly orthogonal directions. The effect is condition-dependent: the WEC and DEC are more sensitive to observer choice than the NEC, because they involve timelike observers (parameterized by rapidity) rather than null directions alone.
For the Alcubierre and Natário metrics, the baseline analysis detects all NEC violation points; Lentz also shows zero missed NEC points but is severely under-resolved ($\ll\!1$ cell; Table 2). Van Den Broeck shows small residual NEC/WEC misses ($0.1\%$$0.4\%$), and SEC misses remain (up to $1.2\%$ for Van Den Broeck). The worst-case WEC severity for boosted observers can be much larger (Section 2.4, Interpretation caveat); the specific ratios are $\zeta_{\max}$-dependent. The violation-set agreement may reflect the high symmetry of these metrics, where the ADM normal is naturally aligned with the principal stress-energy eigenvectors. We hypothesize that the Rodal metric breaks this alignment with its irrotational angular shift component, which would explain both the additional violation points and severity increase; the observer-alignment analysis (Figure 12) confirms this misalignment quantitatively.
Figure 12: Histogram of the alignment angle $\theta=\arccos|\hat{s}\cdot\hat{e}_{\text{spatial}}|$ between the optimizer boost direction $\hat{s}$ (unit spatial 3-vector of the worst-case observer’s velocity) and the spatial part $\hat{e}_{\text{spatial}}$ of the DEC-worst eigenvector of $T^{a}{}_{b}$ (the eigenvector whose principal pressure $|p_{i}|$ most exceeds $\rho$), for the Rodal metric ($50^{3}$ grid, $\zeta_{\max}=5$; see Table 1). Both vectors are normalized in the Euclidean (coordinate) norm; the absolute value folds the eigenvector sign ambiguity into $0^{\circ},90^{\circ}]$. Each panel shows $\sim\!108\,000$ Type I DEC-violating points at the indicated $v_{s}$. Medians range from $76^{\circ}$ ($v_{s}=0.1$) to $85^{\circ}$ ($v_{s}=0.99$); across all four velocities the combined median is $82^{\circ}$ (near-orthogonal), confirming that the optimizer boost direction is far from the eigenvector prediction and that single-frame analysis misses violations because the Eulerian frame does not probe these directions. Only 0.20.21.21.2% of points have alignment below $10^{\circ}$.
Role of algebraic vs. optimization methods.
Since the vast majority of grid points classify as Hawking–Ellis Type I (Table 8), the eigenvalue criteria $\rho\geq 0$, $\rho+p_{i}\geq 0$, and $\rho\geq|p_{i}|$ are the primary determinant of energy condition satisfaction or violation at most points. For Type I, these algebraic slacks provide exact certification: a Type I point satisfies an energy condition if and only if the corresponding eigenvalue inequality holds, regardless of observer, and the result is $\zeta_{\max}$-independent. The observer optimization serves two additional roles: it identifies the worst-case observer (boost direction and rapidity), and it provides a cross-check against the algebraic result. For non-Type-I stress-energy tensors, which do not admit a diagonal eigenvalue representation, the optimization becomes the sole method for evaluating observer-dependent conditions, and the algebraic shortcut does not apply. At such points, the reported WEC/SEC/DEC margins are $\zeta_{\max}$-capped lower bounds: a negative value certifies a genuine violation, but a positive value indicates only that no violation was found within the rapidity cap, not that the condition is satisfied for all observers. We validate this pathway with a null-dust (Type II) benchmark in the test suite (see Section 8).
Implications for warp drive engineering.
Any claim that a warp metric satisfies an energy condition must specify which observers were tested and with what coverage. The original claim that “the Lentz metric has positive energy” was shown by Celmaster and Rubin [21] to be incorrect even for Eulerian observers. More generally, our results for Rodal and WarpShell demonstrate that single-frame analysis can miss substantial violation fractions, while for other metrics the Eulerian frame detects all violation points. Continuous observer optimization, as implemented in warpax, quantifies both the spatial extent of missed violations and the severity increase for non-Eulerian observers.
Comparison with WarpFactory.
WarpFactory [4, 5, 24] samples $\sim\!1000$ observer directions and velocities at each grid point and takes the minimum energy condition value over the sample set. This discrete-sampling approach correctly identifies observer dependence, but its accuracy depends on the sampling density: narrow violation cones in the observer manifold may be missed if no sample falls within them. A Fibonacci-lattice benchmark (Figure 13) confirms this quantitatively: even $10^{4}$ deterministic samples per point saturate at $\sim\!93$% DEC detection for the Rodal metric, with a sampled minimum margin three orders of magnitude looser than the observer-optimized result. warpax complements sampling with continuous observer optimization (Section 2.4), which targets the worst-case observer via gradient descent with multi-start initialization. The two approaches are complementary: WarpFactory provides a well-tested MATLAB implementation for rapid exploration, while warpax provides gradient-based observer optimization, autodiff curvature, and Hawking–Ellis classification in a GPU-accelerated Python ecosystem. Our Eulerian-baseline NEC/WEC results match WarpFactory’s published values for the Alcubierre metric at matched parameters; residual $\sim\!1$% relative differences are consistent with autodiff vs. finite-difference curvature computation.
Computational cost.
The observer optimization is the dominant computational cost. For a single metric at $50^{3}$ resolution with $N_{\text{starts}}=8$ optimizer runs per point, the full analysis takes on the order of minutes on GPU hardware (compared to seconds for the Eulerian-only analysis). For Type I points, the Hawking–Ellis classification provides exact algebraic slacks (for the implemented regularized metric) that bypass the optimization entirely for violation detection; the optimizer cost is then justified primarily by worst-observer identification.
Limitations.
The observer optimizer with $N_{\text{starts}}=8$ random restarts is not guaranteed to find the global minimum; for highly non-convex objectives, some local minima may be missed. The soft rapidity cap $\zeta_{\max}\tanh(|\mathbf{w}|/\zeta_{\max})$ attenuates gradients near the rapidity bound, which could impede finding near-lightlike worst-case observers. Concretely, the $\tanh$ map has derivative $\mathrm{sech}^{2}(|\mathbf{w}|/\zeta_{\max})$, which decays exponentially as $|\mathbf{w}|\to\infty$; an optimizer step near the rapidity ceiling therefore receives a vanishingly small gradient signal, and worst-case observers that require $\zeta$ close to $\zeta_{\max}$ may not be found to full precision. A constrained optimizer (projected gradient or L-BFGS-B with hard bounds) would eliminate this failure mode and is planned for a future release. The boost-vector normalization uses the smooth approximation $|\mathbf{w}|_{\varepsilon}=\sqrt{\mathbf{w}\cdot\mathbf{w}+\varepsilon^{2}}$ (with $\varepsilon=10^{-12}$) to remain $C^{\infty}$ at $\mathbf{w}=0$, avoiding discontinuous gradients that could confuse quasi-Newton Hessian updates near the Eulerian point. The ablation study in Section 7.2 confirms that margins and missed-violation fractions are effectively converged by $N_{\text{starts}}=4$, with $N_{\text{starts}}=8$ providing a safety margin.
Since most warp metrics in our test suite are predominantly Hawking–Ellis Type I, the observer optimizer is not extensively demonstrated on a regime where it is strictly necessary for violation detection (algebraic checks suffice for Type I). To validate the non-Type-I pathway, we include a null-dust benchmark (Type II by construction, $T_{ab}=\Phi^{2}k_{a}k_{b}$) in the test suite, verifying that the classifier correctly identifies the null eigenvector structure and the optimizer recovers the expected NEC saturation ($T_{ab}k^{a}k^{b}=0$ to $10^{-12}$ residual) and non-negative WEC ($T_{ab}u^{a}u^{b}\geq 0$) for all tested observers; full numerical results are in the repository test suite. Full Richardson extrapolation (Section 7) was performed for the Alcubierre metric; the three-point data is non-monotone, so the assumed second-order extrapolation confirms resolution stability rather than a measured convergence rate. Rodal and WarpShell were assessed via resolution stability checks (missed-fraction and margin variation across $25^{3}$, $50^{3}$, and $100^{3}$ grids) but were not subjected to formal Richardson extrapolation, so the reported violation fractions for those metrics carry additional resolution uncertainty.
Comparison with dense sampling.
To validate that observer optimization provides genuine value beyond dense observer sampling, we evaluate detection rates at the Eulerian-missed violation points for each metric. For the Rodal NEC and WEC, dense sampling at $10^{4}$ random directions detects 100% of the violations that the optimizer finds. However, for the Rodal DEC, sampling plateaus at only 54% detection even at $10^{4}$ directions, while in our tests the optimizer recovered 100% of the algebraic Type I violation set, a clear case where the violation cone in observer space is too narrow for random sampling to cover reliably. These results demonstrate that while dense sampling suffices for NEC and WEC in most cases, gradient-based optimization substantially outperforms random sampling for DEC analysis, where the violation cone in observer space can be too narrow for random sampling to cover reliably. For Type I points, the algebraic eigenvalue check provides exact DEC truth (for the implemented regularized metric) without any observer search; the optimization advantage is relevant for analyses that rely on observer sampling. Table 10 summarizes these detection rates.
To further quantify the sampling–optimization gap, we evaluate a deterministic Fibonacci-lattice sampler on the Rodal DEC (Figure 13). Using $N_{\text{dir}}$ directions $\times$ 10 rapidity values (up to $\zeta_{\max}=5$), the detection rate of algebraic DEC violations saturates at $\sim\!93$% by $N_{\text{dir}}=50$ (500 total samples) and does not improve meaningfully at $N_{\text{dir}}=1000$ (10 000 total samples); the remaining $\sim\!7$% of violating points have DEC margins too shallow for any fixed sampling grid to resolve. Moreover, the sampled minimum DEC margin is $1.1\times 10^{-2}$, a relatively deep violation easily captured by any reasonable sampling density. In contrast, the observer optimizer identifies violations with margins as small as $2.0\times 10^{-6}$, three orders of magnitude closer to the satisfaction boundary. These shallow violations, concentrated in narrow observer-space cones, are precisely the points that fixed sampling grids fail to resolve.
Table 10: Percentage of Eulerian-missed violations detected by dense random observer sampling at $10^{4}$ directions, vs. 100% for the observer optimizer ($\zeta_{\max}=5$). “Ground truth” here is the algebraic Type I violation set.
Metric NEC WEC DEC
Rodal 100% 100% 54%
Figure 13: Fibonacci-lattice DEC sampling convergence for the Rodal metric ($25^{3}$ grid, $\zeta_{\max}=5$; see Table 1). Detection rate (left axis, circles) saturates at $\sim\!93$% up to the tested density ($\leq 10^{4}$ directions per point). Dashed line: observer optimizer detection at 100%. The gap reflects violating points with DEC margins too shallow for any fixed sampling grid to resolve.
WarpShell regularity.
The WarpShell implementation uses $C^{2}$-smooth quintic Hermite-smoothstep transitions (Appendix A motivates this choice over $C^{1}$ cubic) for all three metric components:
  • Shell boundaries ($R_{1}=0.5$, $R_{2}=1.0$): centers of the Hermite-smoothstep shell indicator ($C^{2}$ smooth) that controls the spatial-metric conformal factor.
  • Lapse transition: smooth Hermite blending to unity beyond $R_{2}+\text{smooth\_width}$. The interior is only approximately flat (the shell indicator evaluates to $\sim\!0.02$ at $r=0$).
  • Shift transition: smooth Hermite blending to zero beyond $R_{2}$.
All three transitions are $C^{2}$-smooth, with continuous first and second derivatives at all transition boundaries. The $C^{2}$ quintic smoothstep $6t^{5}-15t^{4}+10t^{3}$ satisfies $f(0)=0$, $f(1)=1$, $f^{\prime}(0)=f^{\prime}(1)=0$, and $f^{\prime\prime}(0)=f^{\prime\prime}(1)=0$, ensuring that the Riemann tensor components (which involve second derivatives of the metric) are continuous across transition boundaries. Appendix A shows that replacing the $C^{2}$ quintic with a $C^{1}$ cubic smoothstep does not change the energy condition results; the $C^{2}$ choice is motivated by Riemann continuity.
The spatial-metric conformal factor $\gamma_{rr}^{\text{eff}}-1$ decays exponentially but never reaches zero exactly. This residual tail is the source of curvature at large $r$.
We performed a sensitivity analysis excluding boundary bands of width $\delta=0.05$$0.5$ around the shell boundaries $R_{1}$, $R_{2}$. The WEC missed violations are largely insensitive to boundary exclusion: at $v_{s}=0.1$, fewer than 1% of missed WEC points lie within the boundary band, and the filtered missed fraction remains below $0.1\%$ regardless of $\delta$. In contrast, NEC and DEC missed violations are concentrated at boundaries (100% and $87\%$ respectively at $\delta=0.5$), indicating these are transition-zone artifacts. The SEC missed fraction (<0.1<\!0.1%) is similarly robust to boundary exclusion ($1\%$ at boundaries). The radial profile reveals that WEC violations concentrate at $r\approx 3.5$$4.0$, well outside the lapse transition boundary ($R_{2}+\text{smooth\_width}$). These violations arise from the residual spatial-metric curvature of the smooth $\tanh$ tail, not from transition-zone artifacts. The computed curvature is exact for this regularized numerical profile (via autodiff), so the violations are genuine properties of the implemented (regularized) metric, not the intended idealized thin-shell spacetime. Reported pointwise energy conditions for WarpShell should therefore be understood as diagnostics of this specific regularized implementation; the large stress-energy magnitudes (eigenvalues of $T^{a}{}_{b}$ up to $\sim\!10^{25}$ at $50^{3}$ resolution, increasing with grid refinement) near the shell reflect the sharp gradients of the transition functions rather than distributional thin-shell contributions. A tetrad-normalization diagnostic confirms that observer-contracted margins and eigenvalue magnitudes are mutually consistent (ratio $\sim 0.4$ for NEC contractions; see Section 7.3).
The DEC implementation evaluates the three-term diagnostic defined in Section 2.4 (equation 17), checking flux causality, future-directedness, and non-negative energy density simultaneously for both the Eulerian frame and the observer-optimized worst-case direction.
The Rodal metric uses a $C^{\infty}$ regularization $\sqrt{r^{2}+\varepsilon^{2}}$ with $\varepsilon^{2}=10^{-24}$ at $r=0$ in the angular shape function. A sensitivity analysis varying $\varepsilon^{2}\in\{10^{-24},10^{-18},10^{-12},10^{-6}\}$ confirms that violation counts and margins are stable across this range.
To further characterize the Rodal DEC anomaly, we performed a controlled ablation study with three single-variable parameter sweeps (resolution, $N\in\{25,50,100\}$; regularization, $\varepsilon^{2}\in\{10^{-24},10^{-18},10^{-12},10^{-6}\}$; and wall thickness, $\sigma\in\{0.01,0.03,0.1,0.3\}$) using Alcubierre as a control metric. Controlled experiments indicate that the $\sim\!28.5\%$ DEC miss rate is insensitive to both resolution (0.47 pp variation) and regularization (0.0 pp variation), but depends on the wall-thickness parameter $\sigma$, with 12.43 pp variation from $\sigma=0.01$ ($38.50\%$) to $\sigma=0.3$ ($26.07\%$). This geometry dependence is consistent with the Rodal irrotational shift vector, whose angular component $G(r)$ creates anisotropic pressure distributions that sharpen with decreasing wall thickness. The Alcubierre control shows $0.0\%$ DEC miss across all sweep points, confirming experiment validity. Full sweep data are reported in Appendix B.
Two-tier classification tolerance.
The Hawking–Ellis classifier uses a two-tier imaginary-part check to determine whether eigenvalues of $T^{a}{}_{b}$ are effectively real. Tier 1 (absolute): $|\mathrm{Im}\,\lambda_{i}|<\tau\cdot s$, where $s=\max(|\mathrm{Re}\,\lambda|,1)$ and $\tau=10^{-10}$; this prevents small numerical noise ($|\mathrm{Im}|\sim 10^{-8}$ at $|\mathrm{Re}|\sim 10^{3}$) from causing spurious Type IV. Tier 2 (relative): $|\mathrm{Im}\,\lambda_{i}|<3\times 10^{-3}\cdot\max|\mathrm{Re}\,\lambda|$ (unclamped); this catches split degenerate eigenvalue pairs where eig returns $\{\lambda\pm\varepsilon i\}$ with $|\varepsilon|/|\lambda|\ll 1$ even though $|\varepsilon|\gg\tau\cdot s$. This occurs at large $\|T^{a}{}_{b}\|$ (e.g. WarpShell with $|\lambda|$ up to $\sim\!10^{25}$ at $50^{3}$ resolution). Observer-contracted EC margins $T_{ab}k^{a}k^{b}$ are of the same order as the eigenvalue magnitudes (see the tetrad-normalization diagnostic in Section 7.3); the large absolute values ($\mathcal{O}(10^{30})$$\mathcal{O}(10^{35})$ for WEC/SEC) reflect genuine thin-shell curvatures, not normalization artifacts. An eigenvalue spectrum is treated as real if either tier is satisfied. The same scale factor $s$ is used for the degeneracy check. A sensitivity analysis varying $\tau\in\{10^{-12},10^{-10},10^{-8},10^{-6}\}$ and the relative tolerance in $\{10^{-4},3\times 10^{-3},10^{-2}\}$ confirms that the percentage of Type I points and violation counts are stable across this range for both Alcubierre and Rodal. An independent cross-check of Type IV classifications using extended-precision eigensolvers (e.g. mpmath) is planned for a future study; the tolerance sensitivity analysis above already demonstrates that the Type IV fraction is stable under threshold variation, suggesting the classifications reflect genuine near-degeneracy rather than solver artifacts.
Our analysis is restricted to classical energy conditions. Quantum energy inequalities [25, 26] provide weaker but more physical constraints on the stress-energy, and their systematic evaluation over observers remains an open problem. Additionally, our metrics are parameterized by a finite set of shape parameters ($v_{s}$, $R$, $\sigma$), and the results may not generalize to metrics outside this family.
Future directions.
Several extensions are natural: (i) metric-space optimization, where the shape function $f(r_{s})$ itself is optimized to minimize energy condition violations; (ii) integration of quantum inequality bounds [25]; (iii) application to the broader landscape of warp drive metrics, including the Bobrick–Martire classification [22]; and (iv) extension to non-vacuum backgrounds (e.g., cosmological spacetimes).
9 Conclusion
We have presented warpax, a JAX-based, GPU-accelerated toolkit for observer-robust energy condition verification of warp drive spacetimes, providing continuous gradient-based observer optimization over the observer manifold as an alternative to discrete observer sampling. For Type I stress-energy the algebraic eigenvalue check is exact (cap-independent); for non-Type I points the optimizer provides $\zeta_{\max}$-capped diagnostics. Our main findings are:
  1. 1.
    Single-frame analysis can miss substantial violations. For the Rodal metric, the Eulerian DEC analysis fails to detect violations at over 28% of all grid points, and the WEC figure exceeds 15%. For the WarpShell metric (regularized implementation), missed WEC violations are below $0.1\%$ of the grid.
  2. 2.
    Violation severity exceeds single-frame estimates. For the Alcubierre and Natário metrics, the baseline analysis detects all NEC violation points; Lentz shows zero missed NEC/WEC points (nonzero SEC misses exist) and is severely under-resolved ($\ll\!1$ cell; this result is a lower bound). Van Den Broeck exhibits small residual NEC/WEC misses ($0.1\%$$0.4\%$ at $v_{s}=0.5$), and SEC misses persist (up to $1.2\%$ for Van Den Broeck). The observer-optimized WEC severity at $\zeta_{\max}=5$ is much larger (e.g. Alcubierre WEC: $\sim\!90{,}000\times$; this ratio is $\zeta_{\max}$-dependent and grows as $\sim\!e^{2\zeta_{\max}}$). A rapidity-cap sensitivity study (Table 9) confirms that the robust WEC margin scales as $\sim\!e^{2\zeta_{\max}}$, while the NEC margin and violation-point set are cap-independent.
  3. 3.
    Worst-case observers are physically meaningful. Where the optimized WEC observer differs from the Eulerian frame, it corresponds to a Lorentz-boosted observer with non-trivial rapidity. For the Alcubierre metric, we confirm that the boost direction aligns with the bubble propagation axis, with rapidity increasing at higher $v_{s}$. The Rodal and WarpShell metrics show non-zero optimized rapidity at their missed-violation points, but a full directional analysis for these metrics is deferred to future work.
  4. 4.
    Tidal forces at bubble walls are significant. Geodesic deviation through the Alcubierre bubble shows sharp tidal eigenvalue spikes at the bubble wall, consistent with the expectation that steeper walls increase curvature gradients. The peak blueshift at the bubble wall reaches $7.1\times$ at $v_{s}=0.99$ compared to $v_{s}=0.1$.
  5. 5.
    Autodiff eliminates finite-difference truncation error. Richardson extrapolation (with assumed $p=2$; the three-point data is non-monotone) confirms resolution stability for the minimum NEC margin and integrated violation volume, validating the numerical computation. Remaining numerical error is dominated by floating-point roundoff and metric regularization (e.g. the $\sqrt{r^{2}+\varepsilon^{2}}$ smoothing in the Rodal metric).
The complete analysis, including all figures in this paper, is reproducible from the scripts included in the repository. We encourage the warp drive community to complement observer-sampling approaches with continuous observer optimization for rigorous energy condition verification.
[…]
The author acknowledges support from the VinUni Center for Environmental Intelligence (CEI).
[…]
warpax is freely available under the MIT license at https://github.com/anindex/warpax. All results presented in this paper can be reproduced by running the provided scripts. No external datasets are required.
Appendix A WarpShell transition smoothness ablation
The WarpShell metric requires smooth transition functions to connect the flat interior, curved shell, and asymptotically flat exterior regions. Because the Riemann tensor involves second derivatives of the metric, the minimum smoothness class that guarantees continuous curvature across transition boundaries is $C^{2}$. A natural question is whether the simpler $C^{1}$-smooth cubic Hermite smoothstep ($3t^{2}-2t^{3}$) suffices in practice, or whether the additional smoothness of the $C^{2}$ quintic Hermite smoothstep ($6t^{5}-15t^{4}+10t^{3}$) is needed.
The $C^{2}$ polynomial satisfies six boundary conditions: $f(0)=0$, $f(1)=1$, $f^{\prime}(0)=f^{\prime}(1)=0$, and $f^{\prime\prime}(0)=f^{\prime\prime}(1)=0$. The vanishing second derivatives at the transition endpoints ensure that the metric’s second derivatives, and hence the Riemann tensor components computed via jax.jacfwd, are continuous across all shell boundaries. The $C^{1}$ cubic polynomial satisfies only $f(0)=0$, $f(1)=1$, $f^{\prime}(0)=f^{\prime}(1)=0$; its second derivative is discontinuous at the transition endpoints, introducing kinks in the curvature.
Table 11 compares the two choices across all four tested velocities.
Table 11: WarpShell transition smoothness ablation: $C^{1}$ (cubic) vs $C^{2}$ (quintic) across warp velocities ($50^{3}$ grid, $\zeta_{\max}=5$).
$v_{s}$ Type I (%) Type IV (%) $\min\;m_{\mathrm{NEC}}$ $\max|d^{3}\alpha/dx^{3}|$
C1 C2 C1 C2 C1 C2 C1 C2
0.1 99.6 99.6 0.1 0.1 $2.15\times 10^{30}$ $1.29\times 10^{31}$ $55555.50$ $274444.18$
0.5 99.6 99.6 0.1 0.1 $1.64\times 10^{30}$ $7.26\times 10^{31}$ $55555.50$ $274444.18$
0.9 99.6 99.6 0.1 0.1 $1.55\times 10^{32}$ $1.86\times 10^{32}$ $55555.50$ $274444.18$
0.99 99.6 99.6 0.1 0.1 $2.85\times 10^{31}$ $3.24\times 10^{34}$ $55555.50$ $274444.18$
Figure 14 shows the comparison visually.
Figure 14: WarpShell transition smoothness ablation at four warp velocities ($50^{3}$ grid, $\zeta_{\max}=5$). Left: Hawking–Ellis Type I classification percentage. Right: minimum observer-robust NEC margin (log scale). Both quantities are insensitive to the transition order.
The Hawking–Ellis Type I fraction is identical ($99.6\%$) for both smoothness classes across all four velocities, confirming that the classification is insensitive to the transition order at $50^{3}$ resolution. The minimum NEC margin remains comparable in magnitude: both choices produce margins of $\mathcal{O}(10^{30})$$\mathcal{O}(10^{34})$ depending on $v_{s}$, with the largest relative change at $v_{s}=0.1$ where the $C^{2}$ margin ($1.3\times 10^{31}$) is $\sim\!6\times$ larger than the $C^{1}$ value ($2.2\times 10^{30}$). The third-derivative discontinuity $\max|d^{3}\alpha/dx^{3}|$ increases from $5.6\times 10^{4}$ ($C^{1}$) to $2.7\times 10^{5}$ ($C^{2}$), as expected: the quintic polynomial achieves $f^{\prime\prime}=0$ at the endpoints by concentrating curvature change in the transition interior.
Since the $C^{2}$ choice guarantees Riemann continuity by construction without measurably affecting the energy condition results, we adopt the quintic Hermite smoothstep for all WarpShell analyses in this paper.
Appendix B Rodal DEC ablation study
This appendix reports the full parameter sweep data for the Rodal DEC ablation study summarized in Section 8. Three single-variable sweeps were performed with Alcubierre as a control metric; baseline parameters are $v_{s}=0.5$, $N=50$, $\sigma_{\text{Rodal}}=0.03$, $\varepsilon^{2}=10^{-24}$. Table 12 reports the numerical data and Figure 15 visualizes the three sweeps.
Table 12: Rodal DEC ablation sweep data. Each row varies one parameter while holding the others at baseline. The Alcubierre control confirms $0.0\%$ DEC miss throughout.
Sweep Parameter value Rodal DEC miss (%) Alcubierre DEC miss (%)
Resolution ($N$) $25$ 28.21 0.0
$50$ 28.53 0.0
$100$ 28.68 0.0
Regularization ($\varepsilon^{2}$) $10^{-24}$ 28.53
$10^{-18}$ 28.53
$10^{-12}$ 28.53
$10^{-6}$ 28.53
Wall thickness ($\sigma$) $0.01$ 38.50 0.0
$0.03$ 28.53 0.0
$0.1$ 26.94 0.0
$0.3$ 26.07 0.0
Figure 15: Rodal DEC ablation study: three single-variable parameter sweeps with Alcubierre control (red, dashed). Left: resolution ($N=25$, $50$, $100$). Center: regularization ($\varepsilon^{2}=10^{-24}$ to $10^{-6}$). Right: wall thickness ($\sigma=0.01$ to $0.3$; Alcubierre uses proportionally scaled $\sigma$). The DEC miss rate is insensitive to resolution and regularization but varies by 12.43 pp across wall-thickness values.
The sigma sweep reveals that the Rodal DEC anomaly is both a genuine property of the irrotational geometry and modulated by the wall-thickness parameter: thicker walls (smaller $\sigma$, wider transition region) extend the angular shift gradients over a larger spatial volume, increasing the number of grid points where the DEC violation cone is non-empty. (Rodal and Alcubierre share the same $\tanh$ shape function for $F(r)$, so smaller $\sigma$ means a less steep, broader wall in both metrics.) This behavior is absent in Alcubierre, which lacks the angular shift component, confirming that the anomaly is specific to the Rodal geometry rather than a numerical artifact. Future work could investigate whether alternative shape-function profiles reduce the DEC miss rate without sacrificing the irrotational property.
Appendix C Resolution evidence
Table 2 reports the analytical 10–90% wall width (in coordinate units) and corresponding number of grid cells for each metric at the default $50^{3}$ grid. For the Alcubierre, VdB, and Natário metrics, the wall spans $\sim\!1.4$ cells; Richardson extrapolation for the Alcubierre metric (Section 7) confirms resolution stability of the minimum NEC margin (within $\sim\!1$% at $50^{3}$$100^{3}$), and VdB/Natário are expected to behave similarly given their comparable wall resolution. Lentz is severely under-resolved ($\sim\!0.02$ cells) and its zero-missed-point result should be treated as a lower bound. For the Rodal metric ($\sigma=0.03$, wall width $\approx 73$ coordinate units, $\approx 6$ cells), the missed-fraction stability study (Table 6) shows <2<\!2 percentage-point variation across a $64\times$ increase in grid volume ($25^{3}$ to $100^{3}$), confirming that violation fractions are insensitive to resolution at the tested grid spacing. For the WarpShell metric, the regularized thin-shell transition is concentrated over $\lesssim\!1$ cell; Section 7.3 provides a tetrad-normalization diagnostic showing scale consistency, and Section 7.2 confirms margin convergence with $N_{\text{starts}}$. All WarpShell results should therefore be interpreted as diagnostics of the specific regularized implementation rather than the idealized thin-shell limit.
Appendix D Metric definitions
For reproducibility, we collect the ADM $3+1$ decompositions of all six test metrics. In each case the line element takes the general form
$\mathrm{d}s^{2}=-\alpha^{2}\,\mathrm{d}t^{2}+\gamma_{ij}\bigl(\mathrm{d}x^{i}+\beta^{i}\,\mathrm{d}t\bigr)\bigl(\mathrm{d}x^{j}+\beta^{j}\,\mathrm{d}t\bigr),$ (26)
where $\alpha$ is the lapse, $\beta^{i}$ the shift vector, and $\gamma_{ij}$ the spatial metric. All metrics share the bubble center $x_{s}(t)=v_{s}\,t$ and the comoving distance $r_{s}=\sqrt{(x-x_{s})^{2}+y^{2}+z^{2}}$. The standard $\tanh$-smoothed top-hat shape function is
$f(r_{s})=\frac{\tanh\bigl[\sigma(r_{s}+R)\bigr]-\tanh\bigl[\sigma(r_{s}-R)\bigr]}{2\,\tanh(\sigma R)}\,$, (27)
with $f(0)=1$ and $f(r_{s})\to 0$ for $r_{s}\gg R$.
Alcubierre.
$\alpha=1$, $\gamma_{ij}=\delta_{ij}$, $\beta^{i}=(-v_{s}\,f(r_{s}),\;0,\;0)$. The shape function $f$ is given by equation (27) [1].
Lentz.
$\alpha=1$, $\gamma_{ij}=\delta_{ij}$, $\beta^{i}=(-v_{s}\,f_{\diamond}(d),\;0,\;0)$, where $d=|x-x_{s}|+\sqrt{y^{2}+z^{2}}$ is the L1 (diamond) distance and $f_{\diamond}$ is equation (27) evaluated at $d$ instead of $r_{s}$ [17].
Van Den Broeck.
$\alpha=1$, $\beta^{i}=(-v_{s}\,f(r_{s}),\;0,\;0)$, $\gamma_{ij}=B^{2}(r_{s})\,\delta_{ij}$, where $B(r_{s})=1+\alpha_{\text{vdb}}\,f_{B}(r_{s})$. The inner shape function $f_{B}$ uses a separate radius $\tilde{R}$ and width $\sigma_{B}$ [19]:
$f_{B}(r_{s})=\frac{\tanh\bigl[\sigma_{B}(r_{s}+\tilde{R})\bigr]-\tanh\bigl[\sigma_{B}(r_{s}-\tilde{R})\bigr]}{2\,\tanh(\sigma_{B}\tilde{R})}\,$. (28)
Natário.
$\alpha=1$, $\gamma_{ij}=\delta_{ij}$. The divergence-free shift ($\nabla_{i}\beta^{i}=0$) has three nonzero Cartesian components [18]:
$\beta^{x}$ $=-v_{s}\bigl[2\,n(r_{s})+r_{s}\,n^{\prime}(r_{s})\sin^{2}\!\theta\bigr],$
$\beta^{y}$ $=v_{s}\,n^{\prime}(r_{s})\,\frac{(x-x_{s})\,y}{r_{s}}\,,\qquad\beta^{z}=v_{s}\,n^{\prime}(r_{s})\,\frac{(x-x_{s})\,z}{r_{s}}\,$, (29)
where $n(r_{s})=\tfrac{1}{2}\bigl(1-f(r_{s})\bigr)$, $\sin^{2}\!\theta=(y^{2}+z^{2})/r_{s}^{2}$, and primes denote $\mathrm{d}/\mathrm{d}r_{s}$.
Rodal.
$\alpha=1$, $\gamma_{ij}=\delta_{ij}$. The irrotational (curl-free) shift is specified in a spherical tetrad aligned with the propagation axis ($\theta$ measured from +x+x) [20]:
$\hat{\beta}_{r}=-v_{s}\,F(r_{s})\cos\theta\,,\qquad\hat{\beta}_{\theta}=v_{s}\,G(r_{s})\sin\theta\,$, (30)
where $F=f$ (equation 27) and
$G(r_{s})=1-\frac{2\,r_{s}\sigma\sinh(R\sigma)+\cosh(R\sigma)\bigl[\ln\cosh\sigma(r_{s}{-}R)-\ln\cosh\sigma(r_{s}{+}R)\bigr]}{2\,r_{s}\sigma\sinh(R\sigma)}\,.$ (31)
WarpShell.
The only metric in our suite with non-unit lapse and non-flat spatial metric [23]. Using $C^{2}$-smooth quintic Hermite transition functions $S_{\text{warp}}(r)$ and $S_{\text{shell}}(r)$ (Appendix A):
$\alpha$ $=1-S_{\text{shell}}(r)\Bigl(1-\sqrt{1-r_{\mathrm{S}}/r}\Bigr),$
$\beta^{i}$ $=\bigl(-S_{\text{warp}}(r)\,v_{s},\;0,\;0\bigr),$
$\gamma_{ij}$ $=\delta_{ij}+\bigl(\gamma_{rr}^{\text{eff}}(r)-1\bigr)\,\hat{n}_{i}\hat{n}_{j}\,$, (32)
where $r_{\mathrm{S}}$ denotes the Schwarzschild radius parameter ($r_{\mathrm{S}}=5$ in our runs; see Table 1), $r$ the coordinate distance from the bubble center, $\hat{n}_{i}=x_{i}/r$, and
$\gamma_{rr}^{\text{eff}}(r)=1+S_{\text{shell}}(r)\left(\frac{1}{1-r_{\mathrm{S}}/r}-1\right).$ (33)
The flat interior ($r<R_{1}$) and exterior ($r>R_{2}$) are recovered by $S_{\text{shell}}\to 0$; the Schwarzschild-like shell ($R_{1}<r<R_{2}$) has $S_{\text{shell}}\to 1$.